### v1

# Misaki Kojima    Naoki Nishida

Abstract Framework for All-Path Reachability Analysis toward Safety and Liveness Verification (Full Version)
Misaki Kojima    Naoki Nishida
Abstract
An all-path reachability (APR, for short) predicate is a pair of a source set and a target set, which are subsets of an object set. APR predicates have been defined for abstract reduction systems (ARSs, for short) and then extended to logically constrained term rewrite systems (LCTRSs, for short) as pairs of constrained terms that represent sets of terms modeling configurations, states, etc. An APR predicate is said to be partially (or demonically) valid w.r.t. a rewrite system if every finite maximal reduction sequence of the system starting from any element in the source set includes an element in the target set. Partial validity of APR predicates w.r.t. ARSs is defined by means of two inference rules, which can be considered a proof system to construct (possibly infinite) derivation trees for partial validity. On the other hand, a proof system for LCTRSs consists of four inference rules, and thus there is a gap between the inference rules for ARSs and LCTRSs. In this paper, we revisit the framework for APR analysis and adapt it to verification of not only safety but also liveness properties. To this end, we first reformulate an abstract framework for partial validity w.r.t. ARSs so that there is a one-to-one correspondence between the inference rules for partial validity w.r.t. ARSs and LCTRSs. Secondly, we show how to apply APR analysis to safety verification. Thirdly, to apply APR analysis to liveness verification, we introduce a novel stronger validity of APR predicates, called total validity, which requires not only finite but also infinite execution paths to reach target sets. Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to liveness verification.
keywords:
abstract reduction system, cyclic proof, constrained rewriting, runtime-error verification
1 Introduction
Recently, approaches to program verification by means of logically constrained term rewrite systems (LCTRSs, for short) [KN13frocos] are well investigated [FKN17tocl, WM18, CL18, NW18vstte, KN18eptcs, KNS19ss, CLB23, MNKS23wst, KNM25jlamp]. LCTRSs are useful models of not only functional but also imperative programs, which are evaluated sequentially. For instance, equivalence checking by means of LCTRSs is useful to ensure the correctness of terminating functions (cf. [FKN17tocl]). Since the reduction of rewrite systems is in general non-deterministic, rewrite systems are reasonable models of concurrent programs. A transformation of sequential programs into LCTRSs [FKN17tocl, KN18eptcs] has been extended to concurrent programs [KNM25jlamp]. In addition, a method for runtime-error verification by means of all-path reachability analysis of LCTRSs has been developed for, e.g., race and starvation freedom [KN23jlamp, KN23padl, KN24rp].
An all-path reachability (APR, for short) predicate over an object set $A$ is a pair $P\Rightarrow Q$ of a source set $P$ and a target set $Q$, which are subsets of $A$. The APR predicate is said to be partially valid (or demonically valid [CL18]) w.r.t. a rewrite system $\mathcal{S}$ if every finite execution path—a maximal reduction sequence—of $\mathcal{S}$ starting from any element $s$ in $P$ includes an element in $Q$. Partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{S}$ means that every finite execution from $P$ eventually reaches $Q$. The APR problem w.r.t. $\mathcal{S}$ is a problem to determine whether given APR predicates are partially valid w.r.t. $\mathcal{S}$ or not.
An abstract APR framework, which is a framework for abstract reduction systems (ARS, for short), have first been proposed [CL18]. In the framework, partial validity of APR predicates w.r.t. an ARS $\mathcal{A}=(A,\to_{\mathcal{A}})$ is defined by the two rules of the inference system $\mathsf{DVP}$ [CL18], which are shown in Figure˜1. Here, a set $P$ ($\subseteq A$) is said to be $\mathcal{A}$-runnable if $P\neq\emptyset$ and $P$ does not include any normal form of $\mathcal{A}$. Since the inference system $\mathsf{DVP}$ w.r.t. $\mathcal{A}$ can be used to prove partial validity of APR predicates w.r.t. $\mathcal{A}$, it can be considered a proof system to construct (possibly infinite) proof trees of given APR predicates.
$\mbox{\rm{\rm{Subsumption}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\subseteq Q$}\qquad\mbox{\rm{\rm{Step}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \partial_{\mathcal{A}}(P\setminus Q)\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $(P\setminus Q)$ is $\mathcal{A}$-runnable }$
Figure 1: Inference rules of $\mathsf{DVP}$ [CL18] for partial validity of APR predicates w.r.t. an an ARS $\mathcal{A}=(A,\to_{\mathcal{A}})$, where $\partial_{\mathcal{A}}(P)=\{t^{\prime}\in A\mid\exists t\in P.\ t\to_{\mathcal{A}}t^{\prime}\}$.
The abstract APR framework has been adapted to LCTRSs, and the proof system $\mathsf{DCC}$ for partial validity has been introduced [CL18]. An APR predicate over a signature $\Sigma$ is a pair $\langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle$ of constrained terms $\langle s\,|\,\phi\rangle,\langle t\,|\,\psi\rangle$ over $\Sigma$, which represents sets of configurations, states, etc. To simplify the discussion, we assume that the two constrained terms have no shared variables. Note that $\mathcal{R}$ is an ARS for the set of terms over $\Sigma$, and a constrained term $\langle s\,|\,\phi\rangle$ can be considered sets of ground normalized instances of $s$ by means of substitutions satisfying $\phi$. The proof system $\mathsf{DCC}$ w.r.t. an LCTRS $\mathcal{R}$ consists of the four inference rules shown in Figure˜2. Given a set $G$ of APR predicates over $\Sigma$, if we succeed in constructing proof trees for all APR predicates in $G$ by $\mathsf{DCC}$ w.r.t. $\mathcal{R}$, then all APR predicates in $G$ are partially valid w.r.t. $\mathcal{R}$. Note that the application of Subs is followed by Circ or either Axiom or Der.
$\mbox{\rm{\rm{Axiom}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }\penalty 10000\ \mbox{if $\phi$ is unsatisfiable.}$
$\mbox{\rm{\rm{Subs}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle s\,|\,\phi\land\neg(\exists\vec{x}.\ (s=t\land\phi\land\psi))\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }\penalty 10000\ \mbox{if $\exists\vec{x}.\ (s=t\land\phi\land\psi)$ is satisfiable,}$
where $\{\vec{x}\}=\mathcal{V}ar(t,\psi)\setminus\mathcal{V}ar(s,\phi)$.
$\mbox{\rm{\rm{Der}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle s_{1}\,|\,\phi_{1}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ \,\ldots\penalty 10000\ \,\langle s_{n}\,|\,\phi_{n}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }\penalty 10000\ \mbox{if $\langle s\,|\,\phi\rangle$ is $\mathcal{R}$-runnable, }$
where $\Delta_{\mathcal{R}}(\langle s\,|\,\phi\rangle)=\{\langle s_{i}\,|\,\phi_{i}\rangle\mid 1\leq i\leq n\}$ for some $n>0$, and $\Delta_{\mathcal{R}}(\langle s^{\prime}\,|\,\phi^{\prime}\rangle):=\{\langle s^{\prime\prime}\,|\,\phi^{\prime\prime}\rangle\mid\langle s^{\prime}\,|\,\phi^{\prime}\rangle\mathrel{\leadsto_{\mathcal{R}}}\langle s^{\prime\prime}\,|\,\phi^{\prime\prime}\rangle\}$.
$\mbox{\rm{\rm{Circ}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle t^{\prime}\,|\,\phi\wedge\phi^{\prime\prime}\wedge\psi^{\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\quad\langle s\,|\,\phi\wedge\neg\phi^{\prime\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }$
where $\langle s^{\prime}\,|\,\phi^{\prime}\rangle\Rightarrow\langle t^{\prime}\,|\,\psi^{\prime}\rangle$ is a freshly renamed APR predicate in $G$, $\{\vec{x}\}=\mathcal{V}ar(s^{\prime},\phi^{\prime})$, and $\phi^{\prime\prime}=(\exists\vec{x}.\ (s=s^{\prime})\wedge\phi^{\prime})$.
Figure 2: Inference rules of $\mathsf{DCC}$ [CL18] for partial validity of APR predicates (in $G$) of an LCTRS $\mathcal{R}$.
The two proof systems $\mathsf{DVP}$ and $\mathsf{DCC}$ have no one-to-one correspondence between their inference rules. The main role of Circ in $\mathsf{DCC}$ is the introduction of circularity to proof trees of $\mathsf{DCC}$ so as to make the trees finite as in cyclic proofs [Bro05]. Since the initial role of $\mathsf{DVP}$ is to define partial validity, $\mathsf{DVP}$ does not consider circularity and is used to construct possibly infinite proof trees as in $\mbox{LKID}^{\omega}$ [Bro06phd, Chapter 4]. When we do not use Circ in $\mathsf{DCC}$ and allow to construct infinite proof trees, roughly speaking, rules Axiom, Subs, and Der in $\mathsf{DCC}$ correspond to rules Subsumption and Step in $\mathsf{DVP}$ as follows:
  • The application of Axiom and Der in $\mathsf{DCC}$ is simulated by Subsumption and Step in $\mathsf{DVP}$, respectively;
  • the sequential application of Subs and Axiom in $\mathsf{DCC}$ is simulated by Subsumption in $\mathsf{DVP}$, and vice versa;
  • the sequential application of Subs and Der in $\mathsf{DCC}$ is simulated by Step in $\mathsf{DVP}$, and vice versa.
This gap between the proof systems—the absence of one-to-one correspondence between the inference rules—makes the correctness proof of $\mathsf{DCC}$ non-trivial. If $\mathsf{DCC}$ is clearly an instance of $\mathsf{DVP}$, then the correctness of $\mathsf{DCC}$ is immediate from that of $\mathsf{DVP}$; it suffices to show that rules in $\mathsf{DCC}$ are instances of corresponding rules in $\mathsf{DVP}$. In addition, the abstract framework is sufficiently useful to investigate APR-based approaches to verification of, e.g., safety and liveness properties because concurrent transition systems which can be represented by ARSs are usual models for such properties.
To apply APR analysis to runtime-error verification, a weakened but easily-implementable variant of $\mathsf{DCC}$ for LCTRSs has been proposed for simpler APR predicates [KN23jlamp]. The simplification goes well with the verification of safety properties such as race freedom. On the other hand, it requires the verification of liveness properties such as starvation freedom to introduce some approximations. For example, a counter for the waiting time to acquire a semaphore has been introduced to ensure starvation freedom [KN23padl]: states in which the counter value exceeds an upper limit specified in advance for verification are approximately considered error states.
The APR framework seems to be well suited to verification of liveness properties: We let $Q$ be a set of states to be eventually reached; we let $P_{0}$ be either the set of initial states or a set of intermediate states to reach states in $Q$; then, we attempt to prove partial validity of $P_{0}\Rightarrow Q$; the difference operation $P\setminus Q$ remains states in $P$ that have not yet reached any state in $Q$. A weakness of partial validity of APR predicates is that infinite execution paths do not have to be considered, while it does not matter for verification of safety properties: Roughly speaking, reduction sequences ending with error states of safety properties are finite execution paths, and infinite execution paths can be excluded. However, for general runtime-error verification, not only terminating executions but also non-terminating ones have to be taken into account because a runtime error to be verified may happen in a non-terminating execution. For this problem, in [KN23jlamp], simpler APR predicates have been introduced as mentioned above, and a given LCTRS is modified in order to make all finite prefixes of (possibly infinite) execution paths finite execution paths. Unfortunately, the limitation of APR predicates eliminates the good compatibility that exists between $P\setminus Q$ and liveness verification.
In this paper, we aim to develop an abstract foundation for APR analysis toward runtime-error verification. To this end, we revisit the framework for partial validity of APR analysis and adapt it to verification of not only safety but also liveness properties such as race freedom and starvation freedom. We first reformulate an abstract framework for partial validity of APR predicates w.r.t. ARSs so that there is a one-to-one correspondence between the inference rules for partial validity w.r.t. ARSs and LCTRSs. To be more precise, we reformulate inference rules for partial validity of APR predicates w.r.t. ARSs (Section˜3), and adapt a cyclic-proof system for partial validity w.r.t. LCTRSs to ARSs (Section˜4). Secondly, we show how to apply APR analysis to safety verification (Section˜5). Thirdly, to apply APR analysis to liveness verification, we introduce a novel stronger validity of APR problems, called total validity, which requires not only finite but also infinite execution paths to reach target sets (Section˜6.1). Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to liveness verification (Section˜6.2). All omitted proofs of the claims are provided in both the appendix.
As mentioned above, partial validity takes into account finite execution paths only. To apply the APR framework to runtime-error verification, however, all (i.e., not only finite but also infinite) execution paths have to be taken into account. To this end, we introduce a novel notion of validity, called total validity: An APR predicate $P\Rightarrow Q$ over an object set $A$ is said to be totally valid w.r.t. an ARS $\mathcal{A}$ over $A$ if any (possibly infinite) execution path starting from a state in $P$ eventually reaches a state in $Q$. Note that total validity is stronger than partial validity, and thus partial validity is a necessary condition for total validity. We show that for a partially valid APR predicate with a proof tree in cyclic-proof style, there is no cycle in the graph obtained from the tree if and only if the APR predicate is totally valid w.r.t. $\mathcal{A}$. As a consequence, our sufficient condition for total validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is the existence of an APR proof, the proof graph of which is acyclic.
Since “demonical validity” [CL18] (partial validity in this paper) does not consider infinite execution paths and we consider a stronger validity for all execution paths, we renamed “demonical validity” to “partial validity”, which originates from partial and total correctness of programs. Notice that total correctness of programs requires termination of programs, but total validity does not require termination but requires all (possibly infinite) execution paths reach elements in target sets.
The main contributions of this paper are (1) a reformulation of the APR framework for ARSs, (2) a novel notion of validity—total validity—for APR predicates, and (3) a sufficient condition for total validity for, e.g., APR-based verification of liveness properties.
2 Preliminaries
In this section, we briefly recall all-path reachability of ARSs [CL18]. Familiarity with basic notions and notations on ARSs [BN98, Ohl02] is assumed.
An abstract reduction system (ARS, for short) over an object set $A$ is a pair $(A,\to_{\mathcal{A}})$ such that $\to_{\mathcal{A}}$ is a binary relation over $A$. The reflexive closure of $\to_{\mathcal{A}}$ is denoted by $\to_{\mathcal{A}}^{=}$, the transitive closure of $\to_{\mathcal{A}}$ is denoted by $\to_{\mathcal{A}}^{+}$, and the reflexive and transitive closure of $\to_{\mathcal{A}}$ is denoted by $\to_{\mathcal{A}}^{*}$. We denote the set of normal forms (irreducible elements) of $\mathcal{A}$ by $\mathit{NF\!}_{\mathcal{A}}$. An element $s\in A$ is said to have a normal form if there exists a normal form $b\in\mathit{NF\!}_{\mathcal{A}}$ such that $a\mathrel{\to_{\mathcal{A}}^{*}}b$. In the remainder of the paper, we use an ARS $\mathcal{A}=(A,\to_{\mathcal{A}})$ without notice.
The derivative of a set $P$ ($\subseteq A$) is the set $\partial_{\mathcal{A}}(P)=\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}t\}$. The set $P$ is called runnable w.r.t. $\mathcal{A}$ ($\mathcal{A}$-runnable, for short) if $P\neq\emptyset$ and $P\cap\mathit{NF\!}_{\mathcal{A}}=\emptyset$. By definition, it is clear that $\partial_{\mathcal{A}}(P)\neq\emptyset$ if and only if $P\setminus\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$.
We define execution paths of ARSs.
Definition 2.1 (execution path [CL18]).
Let $\tau$ be a (possibly infinite) reduction sequence $s_{1}\to_{\mathcal{A}}s_{2}\to_{\mathcal{A}}\cdots$ of $\mathcal{A}$. We say that $\tau$ is an execution path (of $\mathcal{A}$) if $\tau$ is maximal, i.e., $\tau$ is either finite ending with an irreducible element or infinite.
Note that execution paths are defined coinductively by using two rules in $\mathsf{DVP}$ w.r.t. $\mathcal{A}$ [CL18] (Figure˜1 in this paper).
Next, we define APR predicates over $A$.
Definition 2.2 (APR predicate [CL18]).
An all-path reachability (APR, for short) predicate over $A$ is a pair $P\Rightarrow Q$ of $P,Q\subseteq A$. [^1] APR predicates in this paper are called reachability property and reachability formulas for ARSs and LCTRSs, respectively, in [CL18], and APR problems in [KN23jlamp, KN23padl, KN23jlamp, Kojima24phd]. Since we deal with “all-path reachability” only, we only use “all-path reachability predicate” to uniform the terminologies. Note that $Q$ is not restricted to non-empty sets.
Let $\mathit{Rules}$ be a finite set of inference rules over a set $\mathcal{O}$ of objects. $\widehat{\mathit{Rules}}$ stands for the functional of $\mathit{Rules}$ (see e.g., [CL18corr, Appendix A.2]): For a set $X$ ($\subseteq\mathcal{O}$), $\widehat{\mathit{Rules}}(X)=\{A\mid(\frac{A_{1}\penalty 10000\ \ldots\penalty 10000\ A_{n}}{A}\penalty 10000\ \mbox{\footnotesize if $B_{1}\penalty 10000\ \ldots\penalty 10000\ B_{m}$})\in\mathit{Rules},$ $A,A_{1},\ldots,A_{n}\in\mathcal{O},$ $B_{1},\ldots,B_{m}$ are side conditions, $A_{1},\ldots,A_{n}\in X$, all $B_{1},\ldots,B_{m}$ holds }\}. We denote the greatest fixed point of $\widehat{\mathit{Rules}}$ by $\nu\widehat{\mathit{Rules}}$, i.e., $\nu\widehat{\mathit{Rules}}$ is the greatest set $Y$ ($\subseteq\mathcal{O}$) such that $Y\supseteq\widehat{\mathit{Rules}}(Y)$.
An APR predicate $P\Rightarrow Q$ over $A$ is said to be partially valid (demonically valid [CL18]) w.r.t. $\mathcal{A}$ if each execution path starting from an element in $P$ either eventually reaches an element in $Q$ or is infinite.
Partial validity is defined coinductively as follows.
Definition 2.3 (partial validity [CL18]).
An APR predicate $P\Rightarrow Q$ over $\mathcal{A}$ is said to be partially valid w.r.t. $\mathcal{A}$, written as $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$, if $(P\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}(\mathcal{A})}$, where $\mathsf{DVP}(\mathcal{A})$ consists of the rules Subsumption and Step in Figure˜1. We write $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ if $P\Rightarrow Q$ is not partially valid w.r.t. $\mathcal{A}$.
Regarding partial validity, APR predicates have the following properties.
Proposition 2.4.
Let $P,P^{\prime},Q,Q^{\prime},R\subseteq A$. Then, all of the following hold:
  1. 1.
    if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ and $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P^{\prime}\Rightarrow Q^{\prime}$, then $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}(P\cup P^{\prime})\Rightarrow(Q\cup Q^{\prime})$,
  2. 2.
    if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}(P\cup P^{\prime})\Rightarrow Q$, then $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ and $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P^{\prime}\Rightarrow Q$,
  3. 3.
    if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$, then $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q^{\prime}$ for any $Q^{\prime}$ such that $Q\subseteq Q^{\prime}\subseteq A$,
  4. 4.
    if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ and $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}Q\Rightarrow R$, then $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow R$, and
  5. 5.
    $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow\emptyset$, if and only if any element in $P$ has no normal form, i.e., $\{t\in\mathit{NF\!}_{\mathcal{A}}\mid\exists s\in P.\ s\mathrel{\to_{\mathcal{A}}^{*}}t\}=\emptyset$.
Proof 2.5.
Trivial by definition.
For an APR predicate $P\Rightarrow Q$, regarding partial validity, Proposition˜2.4 (1 and 2) implies soundness and completeness of splitting $P$ to $n$ sets $P_{1},\ldots,P_{n}$ (i.e., $P=P_{1}\cup\cdots\cup P_{n}$):
  • $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$ if and only if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}(P_{1}\cup\cdots\cup P_{n})\Rightarrow Q$ (Proposition˜2.4-12).
On the other hand, splitting the target set $Q$ to $Q_{1},\ldots,Q_{n}$ (i.e., $Q=Q_{1}\cup\cdots\cup Q_{n}$) is sound but not complete in general:
  • if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$, then, $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow(Q_{1}\cup\cdots\cup Q_{n})$ (Proposition˜2.4-1), but
  • the other direction does not hold in general.
The inference rules in $\mathsf{DVP}(\mathcal{A})$ can be considered a proof system for partial validity of APR predicates, i.e., $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ if and only if there exists a (possibly infinite) proof tree obtained from $P\Rightarrow Q$ by applying the rules of $\mathsf{DVP}(\mathcal{A})$.
Example 2.6.
Consider the ARS $\mathcal{A}_{1}=(\{\mathsf{a},\mathsf{b},\mathsf{c},\mathsf{d}\},\to_{\mathcal{A}_{1}})$ such that ${\to_{\mathcal{A}_{1}}}=\{(\mathsf{a},\mathsf{b}),(\mathsf{a},\mathsf{d}),\linebreak(\mathsf{b},\mathsf{a}),(\mathsf{b},\mathsf{c})\}$. We have that $\mathcal{A}_{1}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$ because $(\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\})\in\nu\widehat{\mathsf{DVP}(\mathcal{A}_{1})}$, i.e., we have the following infinite proof tree:
    $\vdots$  (Step) $\{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Step) $\{\mathsf{a},\mathsf{c}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Step) $\{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Step) $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$
On the other hand, we have that $\mathcal{A}_{1}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$ because the finite execution path $\mathsf{a}\to_{\mathcal{A}_{1}}\mathsf{d}$ does not reach $\mathsf{c}$, and thus $(\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\})\notin\nu\widehat{\mathsf{DVP}(\mathcal{A}_{1})}$. In fact, only Step is applicable, and we obtain the following stuck incomplete proof tree:
$\{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c}\}$  (Step) $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$
3 Reformulation of Inference Rules fo ARSs
In this section, we reformulate the proof system $\mathsf{DVP}$ for partial validity w.r.t. ARSs in [CL18] to establish a one-to-one correspondence between the proof rules for partial validity w.r.t. ARSs and LCTRSs.
By abstracting the inference rules in Figure˜2, we obtain inference rules for ARSs, which have one-to-one correspondence with the rules for LCTRSs.
Definition 3.1 ($\mathsf{DVP}_{+}$).
We define $\mathsf{DVP}_{+}(\mathcal{A})$ consisting of the rules in Figure˜3.
In applying Subs and Der of $\mathsf{DVP}_{+}(\mathcal{A})$ to $P\Rightarrow Q$, we can freely split the results of $P\setminus Q$ and $\partial_{\mathcal{A}}(P)$ to $n$ sets $P_{1},\ldots,P_{n}$, respectively. If $P_{i}=\emptyset$ for Subs or Der, then the generated subgoal $P_{i}\Rightarrow Q$ can immediately be proved by Axiom. To avoid such redundant split, we assume that
  • “if $n>1$ then $\emptyset\subset P_{i}\subseteq P$ for all $1\leq i\leq n$” for Subs, and
  • $P_{i}\neq\emptyset$ for all $1\leq i\leq n$” for Der.
Note that for an $\mathcal{A}$-runnable set $P$, we have that $\partial_{\mathcal{A}}(P)\neq\emptyset$ for Der, because $P\setminus\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$. Since constrained rewrite rules of an LCTRS may overlap, we allow $P_{i}$ and $P_{j}$ ($i\neq j$) to have overlaps, i.e., $P_{i}\cap P_{j}\neq\emptyset$.
$\mbox{\rm{\rm{Axiom}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P=\emptyset$.}\qquad\mbox{\rm{\rm{Subs}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ P_{1}\Rightarrow Q\quad\ldots\quad P_{n}\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\cap Q\neq\emptyset$,}$
where $P\setminus Q=P_{1}\cup\cdots\cup P_{n}$ for some $n>0$.
$\mbox{\rm{\rm{Der}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ P_{1}\Rightarrow Q\quad\ldots\quad P_{n}\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\cap Q=\emptyset$ and $P$ is $\mathcal{A}$-runnable,}$
where $\partial_{\mathcal{A}}(P)=P_{1}\cup\cdots\cup P_{n}$ for some $n>0$.
Figure 3: Rules of $\mathsf{DVP}_{+}(\mathcal{A})$ for proving partial validity of APR predicates of an ARS $\mathcal{A}$.
By definition, it is clear that for each APR predicate $P\Rightarrow Q$, at most one rule in $\mathsf{DVP}_{+}(\mathcal{A})$ is applicable (cf. [KN24jip]).
Proposition 3.2.
Let $P,Q\subseteq A$. Then, at most one rule in $\mathsf{DVP}_{+}(\mathcal{A})$ is applicable to the APR predicate $P\Rightarrow Q$.
Proof 3.3 (Proof (Sketch)).
The side conditions of rules in $\mathsf{DVP}_{+}(\mathcal{A})$ are non-overlapping.
For any ARS, $\mathsf{DVP}$ and $\mathsf{DVP}_{+}$ coincide.
Theorem 3.4.
For any ARS $\mathcal{A}$, $\nu\widehat{\mathsf{DVP}(\mathcal{A})}=\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$.
Proof 3.5 (Proof (Sketch)).
The application of Subsumption in $\mathsf{DVP}(\mathcal{A})$ can be simulated by Subs and Axiom in $\mathsf{DVP}_{+}(\mathcal{A})$ (with $n=1$). The application of Step in $\mathsf{DVP}(\mathcal{A})$ can be simulated by Subs and Der in $\mathsf{DVP}_{+}(\mathcal{A})$ (with $n=1$). Therefore, we have that $\nu\widehat{\mathsf{DVP}(\mathcal{A})}\subseteq\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$.
The application of Axiom in $\mathsf{DVP}_{+}(\mathcal{A})$ can be simulated by Subsumption in $\mathsf{DVP}(\mathcal{A})$. The application of Subs in $\mathsf{DVP}_{+}(\mathcal{A})$ is followed by either Axiom or Der in $\mathsf{DVP}_{+}(\mathcal{A})$. The sequential application of Subs and Axiom can be simulated by Subsumption in $\mathsf{DVP}(\mathcal{A})$. The sequential application of Subs and Der can be simulated by Step in $\mathsf{DVP}(\mathcal{A})$. The application of Der in $\mathsf{DVP}_{+}(\mathcal{A})$—a special case of the successive application of Subs and Der—can be simulated by Step in $\mathsf{DVP}(\mathcal{A})$.
The disproof criterion for LCTRSs [KN24jip] is formulated for ARSs as follows.
Proposition 3.6.
Let $P,Q\subseteq A$. If $P\cap Q=\emptyset$ and $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$, then $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$.
Proof 3.7.
We proceed by contradiction. Assume that $P\cap Q=\emptyset$, $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$, and $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$. Then, $P$ includes a normal form $s$ of $\mathcal{A}$: $s\in P\cap\mathit{NF\!}_{\mathcal{A}}$. Since $P\cap Q=\emptyset$, we have that $s\notin Q$. Since $s$ is a normal form of $\mathcal{A}$, $s$ itself is an execution path of $\mathcal{A}$, which starts from $P$. It follows from $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ that $s\in Q$. This contradicts the fact that $s\notin Q$.
Proposition˜3.6 implies an inference rule for disproof (cf. [KN24jip]).
Definition 3.8 ($\mathsf{DVP}_{\bot}$).
The set $\mathsf{DVP}_{\bot}(\mathcal{A})$ of inference rules consists of the rules in $\mathsf{DVP}_{+}(\mathcal{A})$ and the following rule for disproof:
$\mbox{\rm{\rm{Dis}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \bot\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\cap Q=\emptyset$, $P\neq\emptyset$, and $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$ }$
Note that $\bot$ is used as a special case—the invalid one—of APR predicates.
By definition, it is clear that for each APR predicate $P\Rightarrow Q$, exactly one rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ is applicable (cf. [KN24jip]).
Proposition 3.9.
Let $P,Q\subseteq A$. Then, exactly one rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ is applicable to the APR predicate $P\Rightarrow Q$.
Proof 3.10 (Proof (Sketch)).
The side conditions of rules in $\mathsf{DVP}_{\bot}(\mathcal{A})$ are non-overlapping.
4 Cyclic Proof System for APR Problems of ARSs
In this section, we adapt the cyclic-proof system for partial validity of APR predicates w.r.t. LCTRSs to the abstract APR framework reformulated in Section˜3. To this end, we first revisit rule Circ in Figure˜2, and then define APR pre-proofs for partial validity w.r.t. ARSs.
In proving partial validity of APR predicates by means of the inference rules such as $\mathsf{DVP}_{\bot}(\mathcal{A})$, we would like to construct finite proof trees if possible. In the APR framework, a circularity rule, Circ, has been introduced [CL18, Definition 12]. Roughly speaking, the rule can be considered as a composition of “circularity” in cyclic proofs [Bro05] and “cut” in sequent calculus. Rule Circ in Figure˜2 for LCTRSs is formulated for ARSs as follows:
$\mbox{\rm{\rm{Circ}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ Q^{\prime}\Rightarrow Q\qquad(P\setminus P^{\prime})\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $(P^{\prime}\Rightarrow Q^{\prime})\in G$,}$
where $G$ is a given set of APR predicates over $A$. In practice, we start with a single APR predicate to be proved partially valid, $G$ includes the predicate only; when we apply Der to an APR predicate, we add the predicates to $G$; for soundness, each APR predicate applied Circ has an APR predicate applied Der as an ancestor. [^2] For the soundness condition, it suffices that all APR predicates in $G$ are applied Der.
From the viewpoint of practical use, we do not consider “cut” because it is not so easy to split an APR predicate $P\Rightarrow Q$ to appropriate ones $P\Rightarrow R$ and $R\Rightarrow Q$ for some set $R$. Then, the circularity rule without “cut” is formulated for ARSs as follows:
$\mbox{\rm{\rm{Cyc}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $(P^{\prime}\Rightarrow Q^{\prime})\in G$, $P\subseteq P^{\prime}$, and $Q\supseteq Q^{\prime}$,}$
where $G$ is a given set of APR predicates over $A$. To distinguish the circularity rule without “cut” from Circ, we named the former Cyc, which originates from “cyclic proofs”. The role of Cyc is not only “circularity” but also “generalization” of APR predicates: $P^{\prime}\Rightarrow Q^{\prime}$ is more general than $P\Rightarrow Q$ in the sense that if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P^{\prime}\Rightarrow Q^{\prime}$, then $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ (Proposition˜2.4 12). Viewed in this light, if $P=P^{\prime}$ and $Q=Q^{\prime}$, then Cyc is just “circularity” in cyclic proofs, and otherwise, Cyc generalizes $P\Rightarrow Q$ to $P^{\prime}\Rightarrow Q^{\prime}$, which is included in $G$ to be proved partially valid as, e.g., a more general subgoal for the main goal. Note that “cut” is formulated as follows:
$\mbox{\rm{\rm{Cut}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ P\Rightarrow Q^{\prime}\qquad Q^{\prime}\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }$
As in other proof systems, Cut must be powerful but needs a heuristic which may not be easy to automate. As a first step, we do not consider Cut and leave the introduction of Cut to the proof system as future work.
In [KN24jip], a simpler proof system for partial validity w.r.t. LCTRSs has been formulated in the cyclic-proof style [Bro05]. In the proof system, the circularity rule is not explicitly used, but the bud-companion relationship is used instead in proof trees. We reformulate the cyclic-proof system for partial validity w.r.t. ARSs.
Definition 4.1 (derivation tree of $\mathsf{DVP}(\mathcal{A})$).
An APR derivation tree w.r.t. $\mathcal{A}$ is a finite tree $\mathcal{T}=(V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p})$ such that
  • $V$ is a finite set of nodes,
  • a is a total mapping from $V$ to the set of APR predicates over $A$, which includes $\bot$ as an APR predicate,
  • r is a partial mapping from $V$ to $\mathsf{DVP}_{\bot}(\mathcal{A})$,
  • p is a partial mapping from $V$ to $V^{*}$ (we write $\mbox{\sl p}_{j}(v)$ for the $j$-th component of $\mbox{\sl p}(v)$) which is the $j$-th child of $v$, and
  • for all nodes $v\in V$, $\mbox{\sl p}_{j}(v)$ is defined just in case $\mbox{\sl r}(v)$ is a rule with $k$ premises ($1\leq j\leq k)$, and $\frac{\mbox{\sl a}(\mbox{\sl p}_{1}(v))\penalty 10000\ \ldots\penalty 10000\ \mbox{\sl a}(\mbox{\sl p}_{k}(v))}{\mbox{\sl a}(v)}$ is an instance of rule $\mbox{\sl r}(v)$. [^3] If $k=0$, then $\mbox{\sl p}(v)=\epsilon$.
Note that a node $v\in V$ is a leaf if and only if either $\mbox{\sl p}(v)=\epsilon$ or $\mbox{\sl p}(v)$ is undefined. For a node $v$ with $\mbox{\sl a}(v)=(P\Rightarrow Q)$, $\mbox{\sl a}_{L}(v)$ and $\mbox{\sl a}_{R}(v)$ denote $P$ and $Q$, respectively. A leaf $v$ of $\mathcal{T}$ is said to be closed if $\mbox{\sl p}(v)=\epsilon$ or $\mbox{\sl a}(v)=\bot$. A leaf $v$ of $\mathcal{T}$ is said to be open if it is not closed, i.e., $\mbox{\sl p}(v)\neq\epsilon$ and $\mbox{\sl a}(v)\neq\bot$. The set of open leaves of $\mathcal{T}$ is denoted by $V_{\mathit{open}}$.
Introducing the circularity relationship to APR derivation trees, we define APR pre-proofs.
Definition 4.2 (APR pre-proof).
An APR pre-proof for an APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is a pair $(\mathcal{T},\xi)$ of an APR derivation tree $\mathcal{T}=(V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p})$ (with $v_{0}$ the root node) and a partial mapping $\xi$ from $V_{\mathit{open}}$ to $V\setminus V_{\mathit{open}}$ such that $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$, and for any open leaf $v$, if $\xi(v)$ is defined, then $\xi(v)$ is a node of $\mathcal{T}$ such that $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$ [^4] From the viewpoint of Cyc, the condition $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$” can be relaxed to $P\subseteq P^{\prime}$”, where $\mbox{\sl a}(v)=(P\Rightarrow Q)$ and $\mbox{\sl a}(v^{\prime})=(P^{\prime}\Rightarrow Q)$. However, as a starting point, we do not use the relaxed condition, and leave it as future work. and $\mbox{\sl r}(\xi(v))=\mbox{\rm{\rm{Der}}}$. An open leaf $v$ with $\xi(v)$ defined is called a bud node of $\mathcal{T}$, and the node $\xi(v)$ is called a companion of $v$. We denote the set of bud nodes in $V$ by $V_{\mathit{bud}}$ ($\subseteq V_{\mathit{open}}$). The APR pre-proof $(\mathcal{T},\xi)$ is said to be closed if any leaf $v\in V$ is either closed or a bud node. The APR pre-proof $(\mathcal{T},\xi)$ is said to be open if $\mathcal{T}$ is not closed, i.e., there exists an open leaf that is not a bud.
Note that a companion does not have to be an ancestor of its bud nodes. Unlike the usual definition of companions in cyclic-proofs, we required the additional condition $\mbox{\sl r}(\xi(v))=\mbox{\rm{\rm{Der}}}$” the bud node $v$. The reason why requiring the additional condition will be explained later.
Example 4.3.
Let us consider the ARS $\mathcal{A}_{1}$ in Example˜2.6 again. Let $v_{0},v_{1},\ldots,v_{5}$ nodes such that
  • $\mbox{\sl a}(v_{0})=(\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\})$, $\mbox{\sl r}(v_{0})=\mbox{\rm{\rm{Der}}}$, $\mbox{\sl p}(v_{0})=v_{1}$,
  • $\mbox{\sl a}(v_{1})=(\{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c},\mathsf{d}\})$, $\mbox{\sl r}(v_{1})=\mbox{\rm{\rm{Subs}}}$, $\mbox{\sl p}(v_{1})=v_{2}$,
  • $\mbox{\sl a}(v_{2})=(\{\mathsf{b}\}\Rightarrow\{\mathsf{c},\mathsf{d}\})$, $\mbox{\sl r}(v_{2})=\mbox{\rm{\rm{Der}}}$, $\mbox{\sl p}(v_{2})=v_{3}v_{4}$,
  • $\mbox{\sl a}(v_{3})=(\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\})$, neither $\mbox{\sl r}(v_{3})$ nor $\mbox{\sl p}(v_{3})$ is defined,
  • $\mbox{\sl a}(v_{4})=(\{\mathsf{c}\}\Rightarrow\{\mathsf{c},\mathsf{d}\})$, $\mbox{\sl r}(v_{4})=\mbox{\rm{\rm{Subs}}}$, $\mbox{\sl p}(v_{4})=v_{5}$,
  • $\mbox{\sl a}(v_{5})=(\emptyset\Rightarrow\{\mathsf{c},\mathsf{d}\})$, $\mbox{\sl r}(v_{5})=\mbox{\rm{\rm{Axiom}}}$, $\mbox{\sl p}(v_{5})=\epsilon$, and
  • $\xi(v_{3})=v_{0}$.
Then, $((\{v_{0},v_{1},\ldots,v_{5}\},\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is an APR pre-proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$ w.r.t. $\mathcal{A}_{1}$. The APR pre-proof is visualized in Figure˜4 (a), where $\dagger$} indicates the bud-companion relationship, i.e., $\xi(v_{3})=v_{0}$.
$\dagger$} $v_{3}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$            (Axiom) $v_{5}:\penalty 10000\ \emptyset\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Subs) $v_{4}:\penalty 10000\ \{\mathsf{c}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Der)        $v_{2}:\penalty 10000\ \{\mathsf{b}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Subs)      $v_{1}:\penalty 10000\ \{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Der)      $\dagger$} $v_{0}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}{\dagger}$
${v_{5}}$${v_{4}}$${v_{2}}$${v_{1}}$${v_{0}}$
(a) APR pre-proof (b) the proof graph of (a)
Figure 4: An APR pre-proof and its proof graph for the APR predicate $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$.
The proof graph obtained from a closed APR pre-proof is defined as follows.
Definition 4.4 (proof graph).
The proof graph of a closed APR pre-proof $(\mathcal{T},\xi)$ with $\mathcal{T}=(V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p})$ is a directed graph $(V^{\prime},E)$ obtained from $\mathcal{T}$ by identifying each bud node and its companion:
  • $V^{\prime}=V\setminus V_{\mathit{open}}$, where each node in $V$ is associated with the APR predicate $\mbox{\sl a}(v)$, and
  • $(v,v^{\prime})\in E$ if and only if there exists a node $v^{\prime\prime}\in V$ such that $v^{\prime\prime}$ appears in $\mbox{\sl p}(v)$, if $v^{\prime\prime}$ is a bud node then $v^{\prime}$ is a companion of $v^{\prime\prime}$, and otherwise, $v^{\prime\prime}=v^{\prime}$.
For an edge $(v,v^{\prime})\in E$ and a rule name $X\in\{\mbox{\rm{\rm{Axiom}}},\mbox{\rm{\rm{Subs}}},\mbox{\rm{\rm{Der}}},\mbox{\rm{\rm{Dis}}}\}$, we write $v\to_{\mbox{\scriptsize{\rm{\it X}}}}v^{\prime}$ if $\mbox{\sl r}(v)=\mbox{\it X}$.
Example 4.5.
The proof graph of $((\{v_{0},v_{1},\ldots,v_{5}\},\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ in Example˜4.3 (i.e., Figure˜8 (a)) is illustrated in Figure˜8 (b).
By definition, nodes of proof graphs have the following properties.
Proposition 4.6.
Let $(\mathcal{T},\xi)$ be a closed APR pre-proof, $(V,E)$ the proof graph of $(\mathcal{T},\xi)$, $v,v^{\prime}\in V$, $X\in\{\mbox{\rm{\rm{Axiom}}},\mbox{\rm{\rm{Subs}}},\mbox{\rm{\rm{Der}}},\mbox{\rm{\rm{Dis}}}\}$. Then, all of the following statements hold:
  1. 1.
    if $(v,v^{\prime})\in E$, then $\mbox{\sl a}_{R}(v)=\mbox{\sl a}_{R}(v^{\prime})$,
  2. 2.
    if $v\to_{\mbox{\scriptsize{\rm{Subs}}}}v^{\prime}$, then $(\mbox{\sl a}_{L}(v)\setminus\mbox{\sl a}_{R}(v))\supseteq\mbox{\sl a}_{L}(v^{\prime})$,
  3. 3.
    if $v\to_{\mbox{\scriptsize{\rm{Der}}}}v^{\prime}$, then both of the following hold:
    • for any $s\in\mbox{\sl a}_{L}(v)$, there exist a node $v^{\prime\prime}\in V$ and an element $t\in\mbox{\sl a}_{L}(v^{\prime\prime})$ such that $v\to_{\mbox{\scriptsize{\rm{Der}}}}v^{\prime\prime}$ and $s\to_{\mathcal{A}}t\in\mbox{\sl a}_{L}(v^{\prime\prime})$, and
    • for any $t\in\mbox{\sl a}_{L}(v^{\prime})$, there exists an element $s\in\mbox{\sl a}_{L}(v)$ such that $s\to_{\mathcal{A}}t$,
  4. 4.
    if $v\to_{\mbox{\scriptsize{\rm{Dis}}}}v^{\prime}$, then $\mbox{\sl a}(v^{\prime})=\bot$, and
  5. 5.
    if $\mbox{\sl r}(v)=\mbox{\rm{\rm{Axiom}}}$, then $\mbox{\sl a}_{L}(v)=\emptyset$ and $v$ has no outedge.
Proof 4.7.
Trivial by definition.
Cyclic proofs satisfy the global trace condition [Bro05]. Roughly speaking, the global trace condition requires every infinite trace—a sequence of atomic formulas in premises sets of sequents—to follow an infinite path of the proof graph, which goes through infinitely many edges corresponding to the application of rules of inductive predicates. In our case, such edges correspond to those from nodes with Der. To make proof graphs implicitly satisfy the global trace condition, we require companions to be nodes with Der; the form of APR predicates can be considered the same as that of sequents, and each APR predicate has exactly one constrained term corresponding to an atomic formula of premises of sequents; viewed in this light, traces correspond to paths of proof graphs of closed APR pre-proofs; every cycle goes through companions infinitely often, and thus every infinite trace follows an infinite path that goes through infinitely many nodes with Der.
We now define APR (dis)proofs, which are APR pre-proofs satisfying certain conditions.
Definition 4.8 (APR proof and disproof).
An APR pre-proof $((V,a,r,p),\xi)$ for an APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is called an APR proof if the domain of $\xi$ is $V_{\mathit{open}}$, i.e., all open nodes are bud nodes. The APR pre-proof $((V,a,r,p),\xi)$ is called an APR disproof if there exists a node $v\in V$ such that $\mbox{\sl r}(v)=\mbox{\rm{\rm{Dis}}}$.
Note that a closed APR proof is either an APR proof disproof.
Example 4.9.
The APR pre-proof in Figure˜4 (a) is an APR proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$. The APR pre-proof in Figure˜5 is an APR disproof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$.
   $\bot$  (Dis) $\{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c}\}$  (Der) $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$
Figure 5: An APR pre-proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$, which is an APR disproof
APR proofs and disproofs are sound.
Theorem 4.10.
Let $P\Rightarrow Q$ be an APR predicate over $A$. Then, both of the following statements hold:
  1. 1.
    if there exists an APR proof for $P\Rightarrow Q$, then $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$, and
  2. 2.
    if there exists an APR disproof of $P\Rightarrow Q$, then $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$.
Proof 4.11 (Proof (Sketch)).
The first claim is proved by contradiction, analogously to the proof in [CL18, Theorem 3]. For the second claim, the APR disproof for $P\Rightarrow Q$ has a path from the root to a leaf with $\bot$, which implies a finite execution path that starts from $P$ and does not include any element in $Q$.
When $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, the APR predicate $P\Rightarrow Q$ has either an APR proof or an APR disproof. In addition, if the side conditions of rules in $\mathsf{DVP}_{\bot}(\mathcal{A})$ are decidable, then partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
Theorem 4.12.
Let $P\Rightarrow Q$ be an APR predicate over $A$. Then, both of the following statements hold:
  1. 1.
    If $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, then there exists either an APR proof or an APR disproof for $P\Rightarrow Q$, and
  2. 2.
    if $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, and the emptiness, intersection emptiness, and $\mathcal{A}$-runnability problems for $A$ are decidable, then partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
Proof 4.13 (Proof (Sketch)).
For partial validity of $P\Rightarrow Q$, it suffices to consider APR predicates $P^{\prime}\Rightarrow Q$ such that $P^{\prime}\subseteq\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$. Since $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, there are only finitely many APR predicates to be considered, and thus we can construct a closed APR pre-proof by making generated nodes buds as much as possible. Since the closed APR pre-proof is either an APR proof or disproof, the first claim holds.
For the second claim, since the side conditions of rules in $\mathsf{DVP}_{\bot}(\mathcal{A})$ are decidable by the assumption, we can construct a closed APR pre-proof, which is either an APR proof or disproof. Therefore, the partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
The proof system $\mathsf{DCC}$ for LCTRSs [CL18] and its weakened variant [KN23jlamp] are instances of $\mathsf{DVP}_{\bot}(\mathcal{A})$, and APR pre-proofs w.r.t. ARSs, which provide a concrete method to construct (dis)proofs for APR predicates w.r.t. LCTRSs, are used as a foundation for both systems.
When we simultaneously prove two or more APR predicates, i.e., a set $G$ of APR predicates, to be partially valid, the notion of APR derivation trees and pre-proofs can be extended to forests: We consider a set of APR pre-proofs for APR predicates in $G$; the bud-companion relationship is allowed between nodes in different derivation trees.
5 Reduction of Safety Properties to APR predicates
Informally speaking, safety properties specify that “something bad never happens” [BK08, Section 3.3.2]. As a safety property w.r.t. a system and a set $E$ of error states, we consider the property that any (possibly non-terminating) execution of the system never reaches any error state in $E$. This kind of safety properties is formulated for ARSs as follows:
Instance
an ARS $\mathcal{A}=(A,\to_{\mathcal{A}})$ and sets $P,E$ ($\subseteq A$)
Question
Is there a (possibly infinite) execution path $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots$ such that $s_{0}\in P$ and $s_{i}\in E$ for some $i\geq 0$?
This problem is the same as non-reachability from $P$ to $E$. On the other hand, this problem can be reduced to the APR problem [KN23jlamp]. In this section, we revisit the reduction of the non-reachability problem to partial validity of APR predicates w.r.t. LCTRSs, and then reformulate it for ARSs.
Note that in general, the negation of partial validity of APR predicate $P\Rightarrow E$ is not equivalent to non-reachability from $P$ to $E$:
  • The negation of partial validity of $P\Rightarrow E$ is that there exists a finite execution path starting from a state in $P$, which does not include any error state in $E$, and
  • non-reachability from $P$ to $E$ is that there exists no execution path starting from a state in $P$, which does does not include any error state in $E$.
We first revisit the reduction to APR predicates w.r.t. LCTRSs in [KN24jip]. For an LCTRS $\mathcal{R}$ w.r.t. error states represented by constrained terms $\langle e_{1}\,|\,\psi_{1}\rangle,\ldots,\langle e_{k}\,|\,\psi_{k}\rangle$, the reduction proceeds as follows:
  1. 1.
    Fresh constants $\mathsf{success}$ and $\mathsf{error}$ are introduced to the signature of $\mathcal{R}$;
  2. 2.
    constrained rewrite rules to replace any state to $\mathsf{success}$ are added to $\mathcal{R}$;
  3. 3.
    constrained rewrite rules $e_{i}\to\mathsf{error}\penalty 10000\ [\psi_{i}]$ with $1\leq i\leq k$ are added to $\mathcal{R}$;
  4. 4.
    the non-reachability problem w.r.t. the error states $E$ is reduced to the APR predicate $I\Rightarrow\{\mathsf{success}\}$, where $I$ is the set of ground terms representing initial states of a system modeled by $\mathcal{R}$.
Note that the introduced constants are normal forms of the modified LCTRS. The role of $\mathsf{success}$ is to make all prefixes of any reduction sequence from an initial state a finite execution path ending with $\mathsf{success}$. If an initial execution path reaches an error state represented by some $\langle e_{i}\,|\,\psi_{i}\rangle$, then there exists a finite execution path ending with $\mathsf{error}$.
We now adapt the above approach to ARSs. Usually, error states are irreducible, and thus we consider only irreducible error states. If an error state to be considered is reducible, then, as in [KN23jlamp], we introduce a fresh constant such as $\mathsf{error}$ above, together with constrained rewrite rules from the error states to $\mathsf{error}$, which is considered a dummy error state in the modified system.
Let us reconsider to reduce non-reachability (from $P$ to $E$ w.r.t. an ARS $\mathcal{A}$) to a partial valid APR predicate $P\Rightarrow Q$. [^5] We do not consider to reduce to APR predicates of the form $P^{\prime}\Rightarrow Q$ such that $P^{\prime}\neq P$, because we should consider all execution paths starting from $P$, and thus the APR predicate $P\Rightarrow P^{\prime\prime}$ for some $P^{\prime\prime}\subseteq P^{\prime}$ needs to be partially valid w.r.t. $\mathcal{A}$. When we choose an APR predicate of the form $P\Rightarrow Q$, we do not have to find appropriate sets $P^{\prime},P^{\prime\prime}$. For the reduced partial validity, we do not have to take into account infinite execution paths, because any execution path ending with an error state is a finite execution path. If $Q\cap E\neq\emptyset$, then the reduction is not sound, and thus $Q$ should have no error state in $E$: $Q\cap E=\emptyset$. For the partial validity, all normal forms that are reachable from $P$ and are not error states should be included in $Q$: $\{t\in\mathit{NF\!}_{\mathcal{A}}\setminus E\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}\subseteq Q$. If $Q$ includes a reducible state, then the reduction may not be sound: An error state may be reachable from the reducible state. Thus, $Q$ should be a set of normal forms w.r.t. $\mathcal{A}$. We call such $P\Rightarrow Q$ a safety APR predicate for $E$ w.r.t. $\mathcal{A}$.
Definition 5.1 (safety APR predicate for error states).
Let $E\subseteq\mathit{NF\!}_{\mathcal{A}}$ be a set of error states. An APR predicate $P\Rightarrow Q$ is called a safety predicate for $E$ w.r.t. $\mathcal{A}$ if $Q\cap E=\emptyset$, $\{t\in\mathit{NF\!}_{\mathcal{A}}\setminus E\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}\subseteq Q$, and $Q\subseteq\mathit{NF\!}_{\mathcal{A}}$.
Safety APR predicates can be used for verification of safety properties.
Theorem 5.2.
Let $P\subseteq A$, $Q,E\subseteq\mathit{NF\!}_{\mathcal{A}}$, and $P\Rightarrow Q$ be a safety APR predicate for $E$ w.r.t. $\mathcal{A}$. Then, $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ if and only if there is no finite execution path of $\mathcal{A}$ that starts with an element in $P$ and includes an element in $E$.
Proof 5.3.
Proof (Sketch) Both the only-if and if parts are proved by contradiction.
Example 5.4.
Let us consider the program graph in Figure˜6 for Peterson’s mutual exclusion algorithm [BK08, Example 2.25]. Two processes $P_{0},P_{1}$ share Boolean variables $b_{0},b_{1}$ and an integer variable $x$; The Boolean variables $b_{i}$ indicates that $P_{i}$ wants to enter the critical section $\mathsf{crit}_{i}$; The integer variable $x$ stores the identifier of the process that has priority for the critical section at that time. A state of the asynchronous integer transition systems (AITS, for short) with shared variables [BK08] consisting of $\mathit{PG}_{0}$ and $\mathit{PG}_{1}$ are a tuple of a location of $\mathit{PG}_{0}$, a location of $\mathit{PG}_{1}$, and an assignment from variables $b_{0},b_{1},x$ to values. The initial states of the AITS are $\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle$ with $m\in\{0,1\}$. Let $\mathit{State}_{[}i]$ be the set $\{\mathsf{noncrit}_{i},\mathsf{wait}_{i},\mathsf{crit}_{i}\}$ ($i\in\{0,1\}$), $\mathbb{B}$ the set $\{\mathsf{true},\mathsf{false}\}$, and $\mathcal{S}$ the set of assignments for variables $b_{0},b_{1}:\mbox{\sl{bool}}$ and $x:\mbox{\sl{int}}$: $\mathcal{S}=\{\{b_{0}\mapsto v_{0},b_{1}\mapsto v_{1},x\mapsto m\}\mid v_{0},v_{1}\in\mathbb{B},\penalty 10000\ m\in\mathbb{Z}\}$. We denote a state $\langle\,p_{0},p_{1},\sigma\,\rangle$ by $\langle\,p_{0},p_{1},\sigma(b_{0}),\sigma(b_{1}),\sigma(x)\,\rangle$. The AITS is represented by the following ARS:
$\mathcal{A}_{2}=(\{\langle\,p_{0},p_{1},\sigma\,\rangle\mid p_{0}\in\mathit{State}_{0},\,p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\},\to_{\mathcal{A}_{2}})$
such that
  • $\langle\,\mathsf{noncrit}_{0},p_{1},v_{0},v_{1},m\,\rangle\to_{\mathcal{A}_{2}}\langle\,\mathsf{wait}_{0},p_{1},\mathsf{true},v_{1},1\,\rangle$,
  • $\langle\,\mathsf{wait}_{0},p_{1},v_{0},v_{1},m\,\rangle\to_{\mathcal{A}_{2}}\langle\,\mathsf{crit}_{0},p_{1},\mathsf{true},v_{1},m\,\rangle$ if $m=0$ or $v_{1}=\mathsf{false}$,
  • $\langle\,p_{0},\mathsf{crit}_{1},v_{0},v_{1},m\,\rangle\to_{\mathcal{A}_{2}}\langle\,p_{0},\mathsf{noncrit}_{1},v_{,}\mathsf{false},m\,\rangle$,
  • $\langle\,p_{0},\mathsf{noncrit}_{1},v_{0},v_{1},m\,\rangle\to_{\mathcal{A}_{2}}\langle\,p_{0},\mathsf{wait}_{1},v_{0},\mathsf{true},0\,\rangle$,
  • $\langle\,p_{0},\mathsf{wait}_{1},v_{0},v_{1},m\,\rangle\to_{\mathcal{A}_{2}}\langle\,p_{0},\mathsf{crit}_{1},v_{0},\mathsf{true},m\,\rangle$ if $m=1$ or $v_{0}=\mathsf{false}$, and
  • $\langle\,p_{0},\mathsf{crit}_{1},v_{0},v_{1},m\,\rangle\to_{\mathcal{A}_{2}}\langle\,p_{0},\mathsf{noncrit}_{1},v_{0},\mathsf{false},m\,\rangle$,
where $p_{i}\in\mathit{State}_{i}$ with $i\in\{0,1\}$, $v_{0},v_{1}\in\mathbb{B}$, and $m\in\mathbb{Z}$.
Let us consider the race freedom of mutual exclusion—the two processes do not enter their critical sections simultaneously—for the AITS $\mathcal{A}_{2}$. There is no normal form reachable from an initial state. This property will be examined later, together with race freedom. The error states are $\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle$ with $\sigma\in\mathcal{S}$. Since all the error states are reducible w.r.t. $\mathcal{A}_{2}$, we introduce a fresh element $\mathsf{error}$ as an error state to be verified, and add the reduction from the error states to $\mathsf{error}$: $\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle\to_{\mathcal{A}_{2}}\mathsf{error}$, where $\sigma\in\mathcal{S}$. We let $\mathcal{A}_{2}^{\prime}$ be the modified ARS $(\{\langle\,p_{0},p_{1},\sigma\,\rangle\mid p_{0}\in\mathit{State}_{0},\,p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\},\to_{\mathcal{A}_{2}^{\prime}})$, where ${\to_{\mathcal{A}_{2}^{\prime}}}={\to_{\mathcal{A}_{2}}}\cup\{(\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle,\mathsf{error})\mid\sigma\in\mathcal{S}\}$. The race freedom is reduced to the following APR predicate:
$\{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\emptyset$
We obtain an APR proof illustrated in Figure˜7, and thus, by Theorem 4.10, this APR predicate is partially valid w.r.t. $\mathcal{A}_{2}^{\prime}$. Therefore, by Theorem˜5.2, the AITS defined by the program graphs in Figure˜6 is race free and, by Proposition˜2.4 5, there is no normal form reachable from any initial state.
${\mathit{PG}_{0}:\penalty 10000\ \penalty 10000\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\mathsf{noncrit}_{0}$$\footnotesize{\langle b_{0},x\rangle:=\langle\mathsf{true},1\rangle}$$\mathsf{wait}_{0}$$\footnotesize{x=0\,\lor\,\lnot b_{1}}$$\mathsf{crit}_{0}$$\footnotesize{b_{0}:=\mathsf{false}}$
${\mathit{PG}_{1}:\penalty 10000\ \penalty 10000\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\mathsf{noncrit}_{1}$$\footnotesize{\langle b_{1},x\rangle:=\langle\mathsf{true},0\rangle}$$\mathsf{wait}_{1}$$\footnotesize{x=1\,\lor\,\lnot b_{0}}$$\mathsf{crit}_{1}$$\footnotesize{b_{1}:=\mathsf{false}}$
Figure 6: Program graph $\mathit{PG}_{i}$ ($i=0,1$) for Peterson’s mutual exclusion algorithm.
(8)    (3)  (Der) (9)  (Der)    (4)     (9)  (Der) (5)     (Der)       (2)     (2)  (Der) (10)  (Der) (6)     (10)    (2)   (3)     (11)  (Der)    (7)  (Der)      (3)  (Der) (1) $\{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\emptyset$
  • (2)
    $\{\langle\,\mathsf{wait}_{0},\mathsf{noncrit}_{1},\mathsf{true},\mathsf{false},1\,\rangle\}\Rightarrow\emptyset$
  • (3)
    $\{\langle\,\mathsf{noncrit}_{0},\mathsf{wait}_{1},\mathsf{false},\mathsf{true},0\,\rangle\}\Rightarrow\emptyset$
  • (4)
    $\{\langle\,\mathsf{crit}_{0},\mathsf{noncrit}_{1},\mathsf{true},\mathsf{false},1\,\rangle\}\Rightarrow\emptyset$
  • (5)
    $\{\langle\,\mathsf{wait}_{0},\mathsf{wait}_{1},\mathsf{true},\mathsf{true},0\,\rangle\}\Rightarrow\emptyset$
  • (6)
    $\{\langle\,\mathsf{wait}_{0},\mathsf{wait}_{1},\mathsf{true},\mathsf{true},1\,\rangle\}\Rightarrow\emptyset$
  • (7)
    $\{\langle\,\mathsf{noncrit}_{0},\mathsf{crit}_{1},\mathsf{false},\mathsf{true},0\,\rangle\}\Rightarrow\emptyset$
  • (8)
    $\{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},1\,\rangle\}\Rightarrow\emptyset$
  • (9)
    $\{\langle\,\mathsf{crit}_{0},\mathsf{wait}_{1},\mathsf{true},\mathsf{true},0\,\rangle\}\Rightarrow\emptyset$
  • (10)
    $\{\langle\,\mathsf{wait}_{0},\mathsf{crit}_{1},\mathsf{true},\mathsf{true},1\,\rangle\}\Rightarrow\emptyset$
  • (11)
    $\{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},0\,\rangle\}\Rightarrow\emptyset$
Figure 7: An APR proof for $\{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\emptyset$.
For $P\Rightarrow Q$ being a safety APR predicate for $E\subseteq\mathit{NF\!}_{\mathcal{A}}$, the target set $Q$ is required to satisfy that $\{t\in\mathit{NF\!}_{\mathcal{A}}\setminus E\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}\subseteq Q$, while it suffices to satisfy $\{t\in\mathit{NF\!}_{\mathcal{A}}\setminus E\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}=Q$. It may be difficult for a given rewrite system $\mathcal{A}$ to compute the set $\{t\in\mathit{NF\!}_{\mathcal{A}}\setminus E\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$, and thus we allow to use an over-approximation $Q$ such that $\{t\in\mathit{NF\!}_{\mathcal{A}}\setminus E\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}\subseteq Q$. On the other hand, as in [KN23jlamp], a simpler modification is useful: A fresh element $\mathsf{any}$ is introduced to $A$, and extend $\to_{\mathcal{A}}$ to ${\to_{\mathcal{A}}}\cup\{(s,\mathsf{any})\mid s\in A\setminus E\}$. This modification is not an approximation.
Theorem 5.5.
Let $P\Rightarrow Q$ be a safety APR predicate for $E\subseteq\mathit{NF\!}_{\mathcal{A}}$. Then, $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ if and only if $(A\cup\{\mathsf{any}\},{\to_{\mathcal{A}}}\cup\{(s,\mathsf{any})\mid s\in A\setminus E\})\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow\{\mathsf{any}\}$.
Proof 5.6 (Proof (Sketch)).
Both the only-if and if parts are proved by contradiction.
One may think that non-reachability analysis is more suitable for verification of safety properties. However, APR analysis is essentially very similar to non-reachability analysis. For non-reachability from $P$ to $E$, all states reachable from $P$ are verified, unless an error state reachable from $P$ is detected. On the other hand, during the construction of an APR pre-proof, all states reachable from $P$ are verified, unless a subset $P^{\prime}$ with $P^{\prime}\cap E\neq\emptyset$ is detected.
6 Reduction of Liveness Properties to APR Problems
Informally speaking, liveness properties state that “something good” will happen in the future [BK08, Section 3.4]. As a liveness property w.r.t. a system, a source set $P$, and a set $Q$ for “something good”, we consider the property that any (possibly non-terminating) execution of the system starting from any state in $P$ reaches a state in $Q$. This kind of safety properties is formulated for ARSs as follows:
Instance
an ARS $\mathcal{A}=(A,\to_{\mathcal{A}})$ and sets $P,E$ ($\subseteq A$)
Question
Does any (possibly infinite) execution path $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots$ with $s_{0}\in P$ include an element in $Q$ (i.e., $s_{i}\in Q$ for some $i\geq 0$)?
Unfortunately, this problem cannot be reduced to partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$, while $P\Rightarrow Q$ seems natural for liveness properties. The shortcoming is that partial validity does not take into account any infinite execution path. In this section, we introduce a stronger validity, called total validity, which takes into account all (possibly infinite) execution paths. Then, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to verification of liveness properties.
6.1 Total Validity of APR Predicates w.r.t. ARSs
We first introduce a stronger validity which takes into account all execution paths. Total validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ ensures that every execution path starting from an element in $P$ eventually reaches an element in $Q$.
Definition 6.1 (total validity).
An APR predicate $P\Rightarrow Q$ over $A$ is said to be totally valid w.r.t. $\mathcal{A}$, written as $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$, if every (possibly infinite) execution path $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots$ with $s_{0}\in P$ includes an element in $Q$, i.e., $s_{i}\in Q$ for some $i\geq 0$. We write $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$ if $P\Rightarrow Q$ is not totally valid w.r.t. $\mathcal{A}$.
Example 6.2.
[…]
$(12)\penalty 10000\ \{\langle\,\mathsf{wait}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\sigma\in\mathcal{S},\sigma(b_{0})\,{=}\,\mathsf{true}\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\sigma\in\mathcal{S}\}$
By definition, it is clear that total validity implies partial one. In other words, partial validity is a necessary condition for total one. Thus, to prove total validity, we should first prove partial validity and we usually attempt to construct an APR proof. However, the existence of such an APR proof does not ensure total validity. For this reason, we need an additional condition. To this end, we show a sufficient condition for APR proofs to additionally imply total validity.
6.2 A Criterion for Total Validity of Partial Valid APR Predicates
Let us consider an APR proof and its proof graph for $P\Rightarrow Q$. If the proof graph has no cycle, then there is no infinite execution path that does not include any element in $Q$; if there exists such a path, then either it is impossible to obtain a finite APR proof or the proof graph has a cycle. Recall that APR proofs are finite derivation trees.
Theorem 6.3.
Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be a partially valid APR proof of an APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$, and $(V^{\prime},E)$ the proof graph of $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$. Then, $(V^{\prime},E)$ is acyclic if and only if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.
Proof 6.4 (Proof (Sketch)).
Both the only-if and if parts are proved by contradiction. For the if part, we consider a bud node $v$ and its companion $v^{\prime}$ in a cycle such that $\xi(v)=v^{\prime}$. By the assumption, we have that $\mbox{\sl a}_{L}(v)=\mbox{\sl a}_{L}(v^{\prime})$. It follows from Proposition˜4.6 23 that for any element $s\in\mbox{\sl a}_{L}(v)$, there exists an element $s^{\prime}\in\mbox{\sl a}_{L}(v^{\prime})=\mbox{\sl a}_{L}(v)$ such that $s^{\prime}\to_{\mathcal{A}}^{+}s$ and the reduction has no element in $Q$. Using this claim, we can show an infinite execution path that starts with an element in $P$ and does not include any element in $Q$. This contradicts the assumption that $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.
We call a closed APR pre-proof acyclic if the proof graph of the APR pre-proof is acyclic. As a consequence of Theorem˜6.3, for an APR predicate $P\Rightarrow Q$, the existence of an acyclic APR pre-proof is a sufficient condition for total validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$.
Corollary 6.5.
For an APR predicate $P\Rightarrow Q$ over $A$, if there exists an acyclic APR proof for $P\Rightarrow Q$ w.r.t. $\mathcal{A}$, then $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.
Once we construct an APR proof, the total validity is decidable.
Corollary 6.6.
Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be an APR proof of a partially valid APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$. Then, both of the following statements hold:
  1. 1.
    $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$ if and only if the proof graph of $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is acyclic, and
  2. 2.
    total validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
Proof 6.7.
The first claim follows Theorem˜6.3. It is decidable whether a finite graph is acyclic. Thus, the second claim holds.
Example 6.8.
Let us continue with Example˜6.2. We have an acyclic APR proof for the APR predicate (12). Therefore, by Corollary˜6.5, the APR predicate (12) is totally valid w.r.t. $\mathcal{A}_{2}$, and thus the AITS is starvation free for $P_{0}$.
    (Axiom) (17)  (Subs) (13)        (Axiom) (17)  (Subs) (18)  (Der) (14)†        (Axiom) (17)  (Subs) (21)   (14)†  (Der)     (19)      (Der)     (15)        (Axiom) (17)  (Subs) (20)   (14)†  (Der)     (16)  (Der)                  (12)
  • (13)
    $\{\langle\,\mathsf{crit}_{0},p_{1},\mathsf{true},v_{1},m\,\rangle\mid v_{1}\,{\in}\,\mathbb{B},m\,{\in}\,\{0,1\},m\,{=}\,0\lor v_{1}\,{=}\,\mathsf{false}\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\,{\in}\,\mathit{State}_{1},\sigma\,{\in}\,\mathcal{S}\}$
  • (14)
    $\{\langle\,\mathsf{wait}_{0},\mathsf{wait}_{1},\mathsf{true},\mathsf{true},0\,\rangle\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\}$
  • (15)
    $\{\langle\,\mathsf{wait}_{0},\mathsf{crit}_{1},\mathsf{true},v_{1},1\,\rangle\mid v_{1}\in\mathbb{B}\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\}$
  • (16)
    $\{\langle\,\mathsf{wait}_{0},\mathsf{noncrit}_{1},\mathsf{true},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\}$
  • (17)
    $\emptyset\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\}$
  • (18)
    $\{\langle\,\mathsf{crit}_{0},\mathsf{wait}_{1},\mathsf{true},\mathsf{true},0\,\rangle\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\}$
  • (19)
    $\{\langle\,\mathsf{wait}_{0},\mathsf{noncrit}_{1},\mathsf{true},\mathsf{false},1\,\rangle\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\}$
  • (20)
    $\{\langle\,\mathsf{crit}_{0},\mathsf{noncrit}_{1},\mathsf{true},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\}$
  • (21)
    $\{\langle\,\mathsf{crit}_{0},\mathsf{noncrit}_{1},\mathsf{true},\mathsf{false},1\,\rangle\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\}$
  • (22)
    $\{\langle\,\mathsf{wait}_{0},\mathsf{wait}_{1},\mathsf{true},\mathsf{true},0\,\rangle\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\}$
Figure 8: An APR proof for (12), where $\dagger$} indicates bud-companion relationships.
7 Related Work
The work in this paper is the adaptation of the APR frameworks in [CL18, KN23jlamp] to ARSs, providing an abstract foundation for APR analysis. $\mathbb{K}$ framework [RS10] is a more general setting of rewriting than LCTRSs, and the proof system for all-path reachability has been implemented in $\mathbb{K}$ [SCMMSR14, SCMMSR19]. The race-freedom of Peterson’s mutual exclusion algorithm has been proved in [SCMMSR19] by means of the APR approach, while starvation freedom is not considered. A comparison of all-path reachability logic with CTL* can be seen in [SCMMSR19]. Our previous work [KN23padl, KN24rp] for the verification of starvation freedom by means of APR analysis with LCTRSs reduces starvation freedom to safety APR predicates, and does not use the approach to liveness properties in this paper.
Our previous APR-based approach in [KN23jlamp] to safety verification is to add rules to a given LCTRS so as to make any finite prefix of all (possible infinite) execution paths finite execution paths of the modified LCTRS. On the other hand, as described in Section˜5, it suffices to include in the target set all irreducible states that are not error states, and we formulated the necessary condition as safety APR predicates for sets of error states. However, for the inclusion of all non-error irreducible states in the target set, it is enough to add the reduction from all states to a freshly introduced dummy state that is irreducible w.r.t. the modified ARS. Viewed in this light, in practice, these two approaches are similar, and the approach in this paper is an abstract framework of the previous one for LCTRSs.
The cyclic-proof system for partial validity of APR predicates in this paper is a simplified adaptation of the well-known cyclic-proof sytem $\mbox{CLKID}^{\omega}$ [Bro06phd, Chapter 5]. Compared with sequents, APR predicates are very simple, and our APR proofs always satisfy the global trace condition.
8 Conclusion
In this paper, we first reformulated inference rules for partial validity of APR predicates w.r.t. ARSs, and adapted the cyclic-proof system for partial validity w.r.t. LCTRSs to ARSs. Then, we showed how to apply APR analysis to safety verification. Finally, we introduced total validity of APR predicates w.r.t. ARSs and we showed if there is an acyclic APR proof for an ARP predicate w.r.t. an ARS, then the APR predicate is totally valid w.r.t. the ARS.
Our APR pre-proofs do not consider the generalization for bud nodes. To be more precise, in the current definition, a bud and its companion have the same APR predicate. As described in Footnote˜4, the relaxation of bud-companion relationships—an APR predicate can be more general than that of its companion—is one of our future work.
A future direction of this research is to deal with the full version of the rule Circ, i.e., “cut” in APR pre-proofs. The other future direction is to introduce process-fairness to the abstract APR framework formulated in this paper.
Appendix A Omitted Proofs
See 3.2
Proof A.1.
We make a case analysis depending on whether $P=\emptyset$.
  • Case where $P=\emptyset$. Rule Axiom is applicable, but the others are not.
  • Case where $P\neq\emptyset$. We further make a case analysis depending on whether $P\cap Q=\emptyset$.
    • Case where $P\cap Q=\emptyset$. We make a case analysis depending on whether $P\cap\mathit{NF\!}_{\mathcal{A}}=\emptyset$.
        Case where $P\cap\mathit{NF\!}_{\mathcal{A}}=\emptyset$. Rule Der is applicable, but the other are not.
        Case where $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$. There is no rule applicable to $P\Rightarrow Q$.
    • Case where $P\cap Q\neq\emptyset$. Rule Sub is applicable, but the others are not.
See 3.4
Proof A.2.
We first show that $\nu\widehat{\mathsf{DVP}(\mathcal{A})}\subseteq\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$. The application of Subsumption in $\mathsf{DVP}(\mathcal{A})$ can be simulated by Subs and Axiom in $\mathsf{DVP}_{+}(\mathcal{A})$ (with $n=1$). The application of Step in $\mathsf{DVP}(\mathcal{A})$ can be simulated by Subs and Der in $\mathsf{DVP}_{+}(\mathcal{A})$ (with $n=1$). Therefore, we have that $\nu\widehat{\mathsf{DVP}(\mathcal{A})}\subseteq\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$.
Next, we show that $\nu\widehat{\mathsf{DVP}(\mathcal{A})}\supseteq\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$. The application of Axiom in $\mathsf{DVP}_{+}(\mathcal{A})$ can be simulated by Subsumption in $\mathsf{DVP}(\mathcal{A})$. The application of Subs in $\mathsf{DVP}_{+}(\mathcal{A})$ is followed by either Axiom or Der in $\mathsf{DVP}_{+}(\mathcal{A})$. Let us consider the following case:
$P_{1}\Rightarrow Q$      $P_{n}\Rightarrow Q$  (Subs)      $P\Rightarrow Q$
where $P\cap Q\neq\emptyset$, $P\setminus Q=P_{1}\cup\cdots\cup P_{n}$ for some $n>0$, and if $n>1$ then $P_{i}\neq\emptyset$ for all $1\leq i\leq n$. We make a case analysis depending on whether $P\setminus Q=\emptyset$ or not.
  • Case where $P\setminus Q=\emptyset$. In this case, by definition, we have that $n=1$ and $P_{1}=\emptyset$. Then, by Proposition˜3.2, the only rule Axiom is applicable to $P_{1}\Rightarrow Q$, and we have the following proof:
         (Axiom) $\emptyset\Rightarrow Q$  (Subs) $P\Rightarrow Q$
    Since $P\setminus Q=\emptyset$, we have that $P\subseteq Q$. Therefore, the above successive application of Subs and Axiom can be simulated by Subsumption in $\mathsf{DVP}$:
        (Subsumption) $P\Rightarrow Q$
  • Case where $P\setminus Q\neq\emptyset$. In this case, we have that $P_{i}\neq\emptyset$ for all $1\leq i\leq n$. Since $P_{i}\subseteq P\setminus Q$, we have that $P_{i}\cap Q=\emptyset$. Since $(P\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$, at least one rule is applicable to $P\Rightarrow Q$. By Proposition˜3.2, the applicable rule is Der only. Thus, for each $i\in\{1,\ldots,n\}$, we have the following subtree:
    $P_{i,1}\Rightarrow Q$      $P_{i,m_{i}}\Rightarrow Q$  (Der)       $P_{i}\Rightarrow Q$
    where $\partial_{\mathcal{A}}(P_{i})=P_{i,1}\cup\cdots\cup P_{i,m_{i}}$ for some $m_{i}>0$, and $P_{i,j}\neq\emptyset$ for all $1\leq j\leq m_{i}$. By definition, we have that $P\setminus Q=\bigcup_{i=1}^{n}P_{i}$ and $P_{i}\setminus Q=P_{i}$ for all $1\leq i\leq n$, Hence, we have that $\partial_{\mathcal{A}}(P_{i}\setminus Q)=\partial_{\mathcal{A}}(P_{i})=\bigcup_{j=1}^{m_{i}}P_{i,j}$, and thus $\partial_{\mathcal{A}}(P\setminus Q)=\partial_{\mathcal{A}}(\bigcup_{i=1}^{n}P_{i})=\bigcup_{i=1}^{n}\bigcup_{j=1}^{m_{i}}P_{i,j}$. Thus, we have the following tree of $\mathsf{DVP}(\mathcal{A})$:
    $(\bigcup_{i=1}^{n}\bigcup_{j=1}^{m_{i}}P_{i,j})\Rightarrow Q$  (Step)    $P\Rightarrow Q$
    Since $(P\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$, we have that $(P_{i,j}\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$ for all $1\leq i\leq n$ and $1\leq j\leq m_{i}$. By induction, we have that for all $1\leq i\leq n$, $(P_{i,1}\Rightarrow Q),\ldots,(P_{i,m_{i}}\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}(\mathcal{A})}$, and thus $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P_{i,j}\Rightarrow Q$ for all $1\leq i\leq n$ and $1\leq j\leq m_{i}$. It follows from Proposition˜2.4 that $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}(\bigcup_{i=1}^{n}\bigcup_{j=1}^{m_{i}}P_{i,j})\Rightarrow Q$ and thus $((\bigcup_{i=1}^{n}\bigcup_{j=1}^{m_{i}}P_{i,j})\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}(\mathcal{A})}$. Therefore, we have that $(P\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}(\mathcal{A})}$.
The application of Der in $\mathsf{DVP}_{+}(\mathcal{A})$ is a special case of the successive application of Subs and Der in $\mathsf{DVP}(\mathcal{A})$.
See 3.9
Proof A.3.
By Proposition˜3.2, it suffices to show the case for Dis. When a rule in $\mathsf{DVP}_{+}(\mathcal{A})$ is applicable to $P\Rightarrow Q$, by definition, it is clear that Dis is not applicable to $P\Rightarrow Q$. In the proof of Proposition˜3.2, in the case where no rule is applicable to $P\Rightarrow Q$, we have that $P\neq\emptyset$, $P\cap Q=\emptyset$, and $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$. In this case, by definition, Dis is applicable.
See 4.10
Proof A.4.
We first prove the first claim, which can be proved analogously to the proof in [CL18, Theorem 3]. We proceed by contradiction. Assume that there exists an APR proof of $P\Rightarrow Q$ and $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$. Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be an APR proof for $P\Rightarrow Q$ with root node $v_{0}$, and $(V^{\prime},E^{\prime})$ the proof graph of $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ such that $v_{0}\in V^{\prime}$, where $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$. Since $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$, there exists an execution path $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots\to_{\mathcal{A}}v_{k}$ such that $v_{0}\in P$, $\{v_{0},v_{1},\ldots,v_{n}\}\cap Q=\emptyset$, and $v_{k}\in\mathit{NF\!}_{\mathcal{A}}$. Since $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$, we have that $s_{0}\in\mbox{\sl a}_{L}(v_{0})$. It follows from Proposition˜4.6 that there exist nodes $v_{1},\ldots,v_{k}\in V^{\prime}$ such that $v_{0}\mathrel{(\to_{\mbox{\scriptsize{\rm{Der}}}}\cup\to_{\mbox{\scriptsize{\rm{Subs}}}})}v_{1}\mathrel{(\to_{\mbox{\scriptsize{\rm{Der}}}}\cup\to_{\mbox{\scriptsize{\rm{Subs}}}})}\cdots\mathrel{(\to_{\mbox{\scriptsize{\rm{Der}}}}\cup\to_{\mbox{\scriptsize{\rm{Subs}}}})}v_{k}$. Since $s_{k}\in\mathit{NF\!}_{\mathcal{A}}$, by the construction of proof graphs, $v_{k}$ is a node in $V$ such that $\mbox{\sl r}(v_{k})=\mbox{\rm{\rm{Dis}}}$, and thus $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is an APR disproof. This contradicts the assumption that $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is an APR proof.
Next, we prove the second claim. Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be an APR disproof for $P\Rightarrow Q$ with root node $v_{0}$, where $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$. Then, there exists node $v,v^{\prime}\in V$ such that $\mbox{\sl r}(v)=\mbox{\rm{\rm{Dis}}}$, $\mbox{\sl p}(v)=v^{\prime}$, and $\mbox{\sl a}(v^{\prime})=\bot$. Let $\mbox{\sl a}(v)=(P^{\prime}\Rightarrow Q)$. Then, by the condition of $v,v^{\prime}$, we have that $P^{\prime}\neq\emptyset$, $P^{\prime}\cap Q=\emptyset$, and $P^{\prime}\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$. Thus, there exists an element $s\in P^{\prime}$ such that $s\in\mathit{NF\!}_{\mathcal{A}}$ and $s\notin Q$. By the construction of APR pre-proofs, there exist nodes $v_{1},\ldots,v_{k}$ such that $v_{0}\mathrel{(\to_{\mbox{\scriptsize{\rm{Der}}}}\cup\to_{\mbox{\scriptsize{\rm{Subs}}}})}v_{1}\mathrel{(\to_{\mbox{\scriptsize{\rm{Der}}}}\cup\to_{\mbox{\scriptsize{\rm{Subs}}}})}\cdots\mathrel{(\to_{\mbox{\scriptsize{\rm{Der}}}}\cup\to_{\mbox{\scriptsize{\rm{Subs}}}})}v_{k}\mathrel{(\to_{\mbox{\scriptsize{\rm{Der}}}}\cup\to_{\mbox{\scriptsize{\rm{Subs}}}})}v\mathrel{(\to_{\mbox{\scriptsize{\rm{Der}}}}\cup\to_{\mbox{\scriptsize{\rm{Subs}}}})}v^{\prime}$. It is clear that Proposition˜4.6 holds for paths of derivation trees, and thus, by Proposition˜4.6, there exist elements $s_{0},s_{1},\ldots,s_{k}$ such that $s_{0}\mathrel{\to_{\mathcal{A}}^{=}}s_{1}\mathrel{\to_{\mathcal{A}}^{=}}\cdots\mathrel{\to_{\mathcal{A}}^{=}}s_{k}\mathrel{\to_{\mathcal{A}}^{=}}s$ and $s_{i}\in\mbox{\sl a}_{L}(v_{i})$ and $\mbox{\sl a}_{L}(v_{i})\cap Q=\emptyset$ for all $0\leq i\leq k$. Thus, we have that $s_{i}\notin Q$ for all $0\leq i\leq k$, and thus $s_{0}\mathrel{\to_{\mathcal{A}}^{=}}s_{1}\mathrel{\to_{\mathcal{A}}^{=}}\cdots\mathrel{\to_{\mathcal{A}}^{=}}s_{k}\mathrel{\to_{\mathcal{A}}^{=}}s$ is a finite execution path that does not include any element in $Q$. Therefore, we have that $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$.
See 4.12
Proof A.5.
We first show a naive construction of APR pre-proofs of $P\Rightarrow Q$, where we do not split any source sets. It follows from Proposition˜3.9 that for each APR predicate, there is exactly one applicable rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$. Let $v_{0}$ be the root of a derivation tree we construct now, where $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$. In the following, we use $V$ as a set of nodes that have already been applied a rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ or considered a bud node; we use $V_{\mathrm{Der}}$ as a set of nodes that can be companions of other nodes; we use $U$ as a queue of nodes that have not been applied any rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ yet. Then, we construct an APR pre-proof by breadth-first search (cf. [KN24jip]):
  1. 1.
    $V:=\{v_{0}\}$, $V_{\mathrm{Der}}:=\emptyset$, and insert $v_{0}$ to $U$;
  2. 2.
    we repeat the following until $U$ has no node:
    1. (a)
      take a node $v$ from $U$, and $V:=V\cup\{v\}$;
    2. (b)
      if there exists a node $v^{\prime}\in V_{\mathrm{Der}}$ with $\mbox{\sl a}(v^{\prime})=\mbox{\sl a}(v)$, then $\xi(v):=v^{\prime}$ and go to a.;
    3. (c)
      if Axiom is applicable to $\mbox{\sl a}(v)$, then $\mbox{\sl r}(v):=\mbox{{\rm{Axiom}}}$, $\mbox{\sl p}(v):=\epsilon$;
    4. (d)
      if Subs is applicable to $\mathit{a}(v)$, then create a node $v^{\prime}$, insert $v^{\prime}$ to $U$, $\mbox{\sl r}(v):=\mbox{{\rm{Subs}}}$, $\mbox{\sl a}(v^{\prime}):=((\mbox{\sl a}_{L}(v)\setminus\mathit{a}_{R}(v))\Rightarrow\mbox{\sl a}_{R}(v))$, and $\mbox{\sl p}(v):=v^{\prime}$;
    5. (e)
      if Der is applicable to $\mathit{a}(v)$, then create a node $v^{\prime}$, insert $v^{\prime}$ to $U$, $\mbox{\sl r}(v):=\mbox{{\rm{Der}}}$, $\mbox{\sl a}(v^{\prime}):=(\partial_{\mathcal{A}}(\mbox{\sl a}_{L}(v))\Rightarrow\mbox{\sl a}_{R}(v))$, $\mbox{\sl p}(v):=v^{\prime}$, and $V_{\mathrm{Der}}:=V_{\mathrm{Der}}\cup\{v\}$;
    6. (f)
      if Dis is applicable to $\mathit{a}(v)$, then create a node $v^{\prime}$, $V:=V\cup\{v^{\prime}\}$, $\mbox{\sl r}(v):=\mbox{{\rm{Dis}}}$, $\mbox{\sl a}(v^{\prime}):=\bot$, and $\mbox{\sl p}(v):=v^{\prime}$;
  3. 3.
    output $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$.
It is clear that the above procedure is deterministic and the output is a closed APR pre-proof.
We now prove the first claim. To prove partial validity of $P\Rightarrow Q$, it suffices to consider APR predicates $P^{\prime}\Rightarrow Q$ such that $P^{\prime}\subseteq\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$. Since $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, there are only finitely many APR predicates to be considered. If the above procedure does not halt, then it is clear that there are infinitely many APR predicates to be considered for partial validity of $P\Rightarrow Q$, and this contradicts the finiteness of APR predicates to be considered. If there exists a node $v$ with $\mbox{\sl a}(v)=\bot$, then the output is an APR disproof, and otherwise, it is an APR proof.
Finally, we prove the second claim. If the side conditions of rules in $\mathsf{DVP}_{\bot}(\mathcal{A})$ are decidable, then the procedure above halts for any APR predicate if $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite. Therefore, by the assumption, the partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
See 5.2
Proof A.6.
We first prove the only-if part by contradiction. Assume that $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ and there is a finite execution path $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots\to_{\mathcal{A}}s_{k}$ such that $s_{0}\in P$ and $s_{k}\in E\subseteq\mathit{NF\!}_{\mathcal{A}}$. Since $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ and $Q\subseteq\mathit{NF\!}_{\mathcal{A}}$, we have that $s_{k}\in Q$. Since $P\Rightarrow Q$ is a safety APR predicate for $E$ w.r.t. $\mathcal{A}$, we have that $Q\cap E=\emptyset$, and hence $s_{k}\notin E$. This contradicts the fact that $s_{k}\in E$.
Next, we prove the if part by contradiction. Assume that $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ and there is no execution path of $\mathcal{A}$ that starts with an element in $P$ and includes an element in $E$. Since $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$, there exists a finite execution path $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots\to_{\mathcal{A}}s_{k}\in\mathit{NF\!}_{\mathcal{A}}$ such that $s_{0}\in P$ and $\{s_{0},s_{1},\ldots,s_{k}\}\cap Q=\emptyset$. Since $P\Rightarrow Q$ is a safety predicate for $E$ w.r.t. $\mathcal{A}$, we have that $\{t\in\mathit{NF\!}_{\mathcal{A}}\setminus E\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}\subseteq Q$, and hence $s_{k}\in E$. This contradicts the assumption that there is no execution path of $\mathcal{A}$ that starts with an element in $P$ and includes an element in $E$.
See 5.5
Proof A.7.
Let $\mathcal{A}^{\prime}=(A\cup\{\mathsf{any}\},\to_{\mathcal{A}^{\prime}})$, where ${\to_{\mathcal{A}^{\prime}}}=({\to_{\mathcal{A}}}\cup\{(s,\mathsf{any})\mid s\in A\setminus E\})$.
We first prove the only-if part by contradiction. Assume that $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ and $\mathcal{A}^{\prime}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow\{\mathsf{any}\}$. By the construction of $\mathcal{A}^{\prime}$, we have that $\mathit{NF\!}_{\mathcal{A}^{\prime}}=E\cup\{\mathsf{any}\}$. Since $\mathcal{A}^{\prime}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow\{\mathsf{any}\}$, there exists a finite execution path $s_{0}\to_{\mathcal{A}^{\prime}}s_{1}\to_{\mathcal{A}^{\prime}}\cdots\to_{\mathcal{A}^{\prime}}s_{k}\in E$ such that none of $s_{0},s_{1},\ldots,s_{k}$ is $\mathsf{any}$. By the construction of $\mathcal{A}^{\prime}$, we have that $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots\to_{\mathcal{A}}s_{k}\in E$. It follows from Theorem˜5.2 that $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$. This contradicts the assumption that $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$.
Next, we prove the if part by contradiction. Assume that $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ and $\mathcal{A}^{\prime}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow\{\mathsf{any}\}$. Since $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$, by Theorem˜5.2, there exists a finite execution path $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots\to_{\mathcal{A}}s_{k}\in E$. By the construction of $\mathcal{A}^{\prime}$, we have that $s_{0}\to_{\mathcal{A}^{\prime}}s_{1}\to_{\mathcal{A}^{\prime}}\cdots\to_{\mathcal{A}^{\prime}}s_{k}\in E$, and thus $\mathcal{A}^{\prime}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow\{\mathsf{any}\}$ This contradicts the assumption that $\mathcal{A}^{\prime}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow\{\mathsf{any}\}$.
See 6.3
Proof A.8.
We first prove the only-if part by contradiction. Assume that $(V^{\prime},E)$ is acyclic and $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$. Then, there exists a (possibly infinite) execution path $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots$ such that $s_{0}\in P$ and none of $s_{0},s_{1},\ldots$ is in $Q$. We make a case analysis depending on whether the path is finite.
  • Case where the execution path is finite. Let the path have $n$ steps, i.e.., $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots\to_{\mathcal{A}}s_{n}\in\mathit{NF\!}_{\mathcal{A}}$. Since $s_{n}\in\mathit{NF\!}_{\mathcal{A}}\setminus Q$, there exists a node $v\in V$ such that $s_{n}\in\mbox{\sl a}_{L}(v)\subseteq A$, and thus, $\mbox{\sl r}(v)={\rm\textsc{Dis}}$. This contradicts the assumption that $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is an APR proof.
  • Case where the execution path is infinite. Since none of $s_{0},s_{1},\ldots$ is in $Q$, there exists an infinite path of the proof graph $(V^{\prime},E)$. Since $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is an APR proof, $(V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p})$ is a finite derivation tree, and thus $(V^{\prime},E)$ is a finite graph having an infinite path. Thus, $(V^{\prime},E)$ has a cycle. This contradicts the assumption that $(V^{\prime},E)$ is acyclic.
Next, we prove the if part by contradiction. Assume that $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$ and $(V^{\prime},E)$ has a cycle. Then, there exists a bud node $v$ and its companion $v^{\prime}$ in the cycle such that $\xi(v)=v^{\prime}$. By the assumption, we have that $\mbox{\sl a}(v)=\mbox{\sl a}(v^{\prime})$, and thus $\mbox{\sl a}_{L}(v)=\mbox{\sl a}_{L}(v^{\prime})$. Since $v,v^{\prime}$ are in the cycle, there exists a finite path from $v^{\prime}$ to $v$: $v^{\prime}\mathrel{\to_{\mbox{\scriptsize{\rm{Der}}}}}v_{1}\mathrel{({\to_{\mbox{\scriptsize{\rm{Der}}}}}\cup{\to_{\mbox{\scriptsize{\rm{Subs}}}}})^{*}}v$. It follows from Proposition˜4.6 23 that for any element $s\in\mbox{\sl a}_{L}(v)$, there exists an element $s^{\prime}\in\mbox{\sl a}_{L}(v^{\prime})$ such that $s^{\prime}\to_{\mathcal{A}}^{+}s$ and the reduction has no element in $Q$. It follows from $\mbox{\sl a}_{L}(v)=\mbox{\sl a}_{L}(v^{\prime})$ that for any element $s\in\mbox{\sl a}_{L}(v)$, there exists an element $s^{\prime}\in\mbox{\sl a}_{L}(v)$ such that $s^{\prime}\to_{\mathcal{A}}^{+}s$ and the reduction has no element in $Q$. Thus, there exists an infinite execution path that starts with an element $s^{\prime\prime}\in\mbox{\sl a}_{L}(v)$ and does not include any element in $Q$. In the same way, we have a reduction sequence $s_{0}\to_{\mathcal{A}}^{*}s^{\prime\prime}$ such that $s_{0}\in P$, and thus there exists an infinite execution path that starts with an element $s^{\prime\prime}\in\mbox{\sl a}_{L}(v)$ and does not include any element in $Q$. This contradicts the assumption that $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.

### v2

@@
 Abstract
-An all-path reachability (APR, for short) predicate is a pair of a source set and a target set, which are subsets of an object set.
+An all-path reachability (APR, for short) predicate over an object set is a pair of a source set and a target set, which are subsets of the object set.
 APR predicates have been defined for abstract reduction systems (ARSs, for short) and then extended to logically constrained term rewrite systems (LCTRSs, for short) as pairs of constrained terms that represent sets of terms modeling configurations, states, etc.
@@
 ARSs is defined by means of two inference rules, which can be considered a proof system to construct (possibly infinite) derivation trees for partial validity.
-On the other hand, a proof system for LCTRSs consists of four inference rules, and thus there is a gap between the inference rules for ARSs and LCTRSs.
+On the other hand, a proof system for LCTRSs consists of four inference rules, leaving a gap between the inference rules for ARSs and LCTRSs.
 In this paper, we revisit the framework for APR analysis and adapt it to verification of not only safety but also liveness properties.
@@
 Thirdly, to apply APR analysis to liveness verification, we introduce a novel stronger validity of APR predicates, called total validity, which requires not only finite but also infinite execution paths to reach target sets.
-Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to liveness verification.
+Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity.
+The condition implies that if there exists a cyclic-proof tree for an APR predicate, the proof graph of which is acyclic, then the APR predicate is totally valid.
 keywords:
-abstract reduction system, cyclic proof, constrained rewriting, runtime-error verification
+abstract reduction system, reachability, cyclic proof, runtime-error verification
 1 Introduction
 Recently, approaches to program verification by means of logically constrained term rewrite systems (LCTRSs, for short) [KN13frocos] are well investigated [FKN17tocl, WM18, CL18, NW18vstte, KN18eptcs, KNS19ss, CLB23, MNKS23wst, KNM25jlamp].
-LCTRSs are useful models of not only functional but also imperative programs, which are evaluated sequentially.
+LCTRSs are useful models of not only functional but also imperative programs.
 For instance, equivalence checking by means of LCTRSs is useful to ensure the correctness of terminating functions (cf.
@@
 The APR predicate is said to be partially valid (or demonically valid [CL18]) w.r.t.
-a rewrite system $\mathcal{S}$ if every finite execution path—a maximal reduction sequence—of $\mathcal{S}$ starting from any element $s$ in $P$ includes an element in $Q$.
+a rewrite system $\mathcal{A}$ if every finite execution path—a maximal reduction sequence—of $\mathcal{A}$ starting from any element $s$ in $P$ includes an element in $Q$.
 Partial validity of $P\Rightarrow Q$ w.r.t.
-$\mathcal{S}$ means that every finite execution from $P$ eventually reaches $Q$.
+$\mathcal{A}$ means that every finite execution from $P$ eventually reaches $Q$.
 The APR problem w.r.t.
-$\mathcal{S}$ is a problem to determine whether given APR predicates are partially valid w.r.t.
-$\mathcal{S}$ or not.
+$\mathcal{A}$ is a problem to determine whether given APR predicates are partially valid w.r.t.
+$\mathcal{A}$ or not.
 An abstract APR framework, which is a framework for abstract reduction systems (ARS, for short), have first been proposed [CL18].
@@
 $\mathcal{A}$, it can be considered a proof system to construct (possibly infinite) proof trees of given APR predicates.
-$\mbox{\rm{\rm{Subsumption}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\subseteq Q$}\qquad\mbox{\rm{\rm{Step}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \partial_{\mathcal{A}}(P\setminus Q)\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $(P\setminus Q)$ is $\mathcal{A}$-runnable }$
+$\mbox{\rm{\rm{Subsumption}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\subseteq Q$.}\qquad\mbox{\rm{\rm{Step}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \partial_{\mathcal{A}}(P\setminus Q)\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $(P\setminus Q)$ is $\mathcal{A}$-runnable.
+}$
 Figure 1: Inference rules of $\mathsf{DVP}$ [CL18] for partial validity of APR predicates w.r.t.
-an an ARS $\mathcal{A}=(A,\to_{\mathcal{A}})$, where $\partial_{\mathcal{A}}(P)=\{t^{\prime}\in A\mid\exists t\in P.\ t\to_{\mathcal{A}}t^{\prime}\}$.
-The abstract APR framework has been adapted to LCTRSs, and the proof system $\mathsf{DCC}$ for partial validity has been introduced [CL18].
+an ARS $\mathcal{A}=(A,\to_{\mathcal{A}})$, where $\partial_{\mathcal{A}}(P)=\{t^{\prime}\in A\mid\exists t\in P.\ t\to_{\mathcal{A}}t^{\prime}\}$.
+The abstract APR framework has been adapted to LCTRSs, and the proof system $\mathsf{DCC}$ for partial validity has been presented [CL18].
 An APR predicate over a signature $\Sigma$ is a pair $\langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle$ of constrained terms $\langle s\,|\,\phi\rangle,\langle t\,|\,\psi\rangle$ over $\Sigma$, which represents sets of configurations, states, etc.
@@
 where $\{\vec{x}\}=\mathcal{V}ar(t,\psi)\setminus\mathcal{V}ar(s,\phi)$.
-$\mbox{\rm{\rm{Der}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle s_{1}\,|\,\phi_{1}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ \,\ldots\penalty 10000\ \,\langle s_{n}\,|\,\phi_{n}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }\penalty 10000\ \mbox{if $\langle s\,|\,\phi\rangle$ is $\mathcal{R}$-runnable, }$
-where $\Delta_{\mathcal{R}}(\langle s\,|\,\phi\rangle)=\{\langle s_{i}\,|\,\phi_{i}\rangle\mid 1\leq i\leq n\}$ for some $n>0$, and $\Delta_{\mathcal{R}}(\langle s^{\prime}\,|\,\phi^{\prime}\rangle):=\{\langle s^{\prime\prime}\,|\,\phi^{\prime\prime}\rangle\mid\langle s^{\prime}\,|\,\phi^{\prime}\rangle\mathrel{\leadsto_{\mathcal{R}}}\langle s^{\prime\prime}\,|\,\phi^{\prime\prime}\rangle\}$.
-$\mbox{\rm{\rm{Circ}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle t^{\prime}\,|\,\phi\wedge\phi^{\prime\prime}\wedge\psi^{\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\quad\langle s\,|\,\phi\wedge\neg\phi^{\prime\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }$
-where $\langle s^{\prime}\,|\,\phi^{\prime}\rangle\Rightarrow\langle t^{\prime}\,|\,\psi^{\prime}\rangle$ is a freshly renamed APR predicate in $G$, $\{\vec{x}\}=\mathcal{V}ar(s^{\prime},\phi^{\prime})$, and $\phi^{\prime\prime}=(\exists\vec{x}.\ (s=s^{\prime})\wedge\phi^{\prime})$.
+$\mbox{\rm{\rm{Der}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle s_{1}\,|\,\phi_{1}\rangle\Rightarrow\langle t\,|\,\psi\rangle\quad\ldots\quad\langle s_{n}\,|\,\phi_{n}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }\penalty 10000\ \mbox{if $\langle s\,|\,\phi\rangle$ is $\mathcal{R}$-runnable, }$
+where $\{\langle s_{i}\,|\,\phi_{i}\rangle\mid 1\leq i\leq n\}=\{\langle s^{\prime}\,|\,\phi^{\prime}\rangle\mid\langle s\,|\,\phi\rangle\mathrel{\leadsto_{\mathcal{R}}}\langle s^{\prime}\,|\,\phi^{\prime}\rangle\}$ for some $n>0$.
+$\mbox{\rm{\rm{Circ}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \langle t^{\prime}\,|\,\phi\wedge\phi^{\prime\prime}\wedge\psi^{\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\qquad\langle s\,|\,\phi\wedge\neg\phi^{\prime\prime}\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }{\penalty 10000\ \langle s\,|\,\phi\rangle\Rightarrow\langle t\,|\,\psi\rangle\penalty 10000\ }$
+where $(\langle s^{\prime}\,|\,\phi^{\prime}\rangle\Rightarrow\langle t^{\prime}\,|\,\psi^{\prime}\rangle)\in G$, $\{\vec{x}\}=\mathcal{V}ar(s^{\prime},\phi^{\prime})$, and $\phi^{\prime\prime}=(\exists\vec{x}.\ (s=s^{\prime})\wedge\phi^{\prime})$.
 Figure 2: Inference rules of $\mathsf{DCC}$ [CL18] for partial validity of APR predicates (in $G$) of an LCTRS $\mathcal{R}$.
@@
 Since the initial role of $\mathsf{DVP}$ is to define partial validity, $\mathsf{DVP}$ does not consider circularity and is used to construct possibly infinite proof trees as in $\mbox{LKID}^{\omega}$ [Bro06phd, Chapter 4].
-When we do not use Circ in $\mathsf{DCC}$ and allow to construct infinite proof trees, roughly speaking, rules Axiom, Subs, and Der in $\mathsf{DCC}$ correspond to rules Subsumption and Step in $\mathsf{DVP}$ as follows:
-  • The application of Axiom and Der in $\mathsf{DCC}$ is simulated by Subsumption and Step in $\mathsf{DVP}$, respectively;
-  • the sequential application of Subs and Axiom in $\mathsf{DCC}$ is simulated by Subsumption in $\mathsf{DVP}$, and vice versa;
-  • the sequential application of Subs and Der in $\mathsf{DCC}$ is simulated by Step in $\mathsf{DVP}$, and vice versa.
-This gap between the proof systems—the absence of one-to-one correspondence between the inference rules—makes the correctness proof of $\mathsf{DCC}$ non-trivial.
+When we do not use Circ in $\mathsf{DCC}$ and allow to construct infinite proof trees, roughly speaking, rules Axiom, Subs, and Der in $\mathsf{DCC}$ correspond to rules Subsumption and Step in $\mathsf{DVP}$, while there is no one-to-one correspondence between the inference rules in $\mathsf{DCC}$ and $\mathsf{DVP}$.
+The absence of one-to-one correspondence makes the correctness proof of $\mathsf{DCC}$ non-trivial.
 If $\mathsf{DCC}$ is clearly an instance of $\mathsf{DVP}$, then the correctness of $\mathsf{DCC}$ is immediate from that of $\mathsf{DVP}$; it suffices to show that rules in $\mathsf{DCC}$ are instances of corresponding rules in $\mathsf{DVP}$.
-In addition, the abstract framework is sufficiently useful to investigate APR-based approaches to verification of, e.g., safety and liveness properties because concurrent transition systems which can be represented by ARSs are usual models for such properties.
+In addition, the abstract framework is sufficiently useful to investigate APR-based approaches to verification of, e.g., safety and liveness properties, because concurrent transition systems which can be represented by ARSs are usual models for such properties.
+In this paper, we aim to develop an abstract foundation for APR analysis toward runtime-error verification.
+To this end, we revisit the framework for partial validity of APR analysis and adapt it to verification of not only safety but also liveness properties such as race and starvation freedom.
 To apply APR analysis to runtime-error verification, a weakened but easily-implementable variant of $\mathsf{DCC}$ for LCTRSs has been proposed for simpler APR predicates [KN23jlamp].
@@
 For example, a counter for the waiting time to acquire a semaphore has been introduced to ensure starvation freedom [KN23padl]: states in which the counter value exceeds an upper limit specified in advance for verification are approximately considered error states.
-The APR framework seems to be well suited to verification of liveness properties: We let $Q$ be a set of states to be eventually reached; we let $P_{0}$ be either the set of initial states or a set of intermediate states to reach states in $Q$; then, we attempt to prove partial validity of $P_{0}\Rightarrow Q$; the difference operation $P\setminus Q$ remains states in $P$ that have not yet reached any state in $Q$.
-A weakness of partial validity of APR predicates is that infinite execution paths do not have to be considered, while it does not matter for verification of safety properties: Roughly speaking, reduction sequences ending with error states of safety properties are finite execution paths, and infinite execution paths can be excluded.
+The APR framework seems to be well suited to verification of liveness properties: We let $P_{0}$ be either the set of initial states or a set of intermediate states to reach states in a target set $Q$; then, we attempt to prove partial validity of $P_{0}\Rightarrow Q$; the difference operation $P\setminus Q$ remains states in $P$ that have not yet reached any state in $Q$.
+The shortcoming of partial validity is that infinite execution paths do not have to be considered, while it does not matter for verification of safety properties: Roughly speaking, reduction sequences ending with error states of safety properties are finite execution paths, and infinite execution paths can be excluded.
 However, for general runtime-error verification, not only terminating executions but also non-terminating ones have to be taken into account because a runtime error to be verified may happen in a non-terminating execution.
-For this problem, in [KN23jlamp], simpler APR predicates have been introduced as mentioned above, and a given LCTRS is modified in order to make all finite prefixes of (possibly infinite) execution paths finite execution paths.
+For this problem, in [KN23jlamp], simpler APR predicates have been introduced as mentioned above, and a given LCTRS is modified in order to make all finite prefixes of (possibly infinite) execution paths finite execution paths of the modified LCTRS.
 Unfortunately, the limitation of APR predicates eliminates the good compatibility that exists between $P\setminus Q$ and liveness verification.
-In this paper, we aim to develop an abstract foundation for APR analysis toward runtime-error verification.
-To this end, we revisit the framework for partial validity of APR analysis and adapt it to verification of not only safety but also liveness properties such as race freedom and starvation freedom.
 We first reformulate an abstract framework for partial validity of APR predicates w.r.t.
@@
 ARSs and LCTRSs.
-To be more precise, we reformulate inference rules for partial validity of APR predicates w.r.t.
+To be more precise, we reformulate inference rules for partial validity w.r.t.
 ARSs (Section˜3), and adapt a cyclic-proof system for partial validity w.r.t.
-LCTRSs to ARSs (Section˜4).
+LCTRSs in [KN24jip, Section 3.2] to ARSs (Section˜4).
 Secondly, we show how to apply APR analysis to safety verification (Section˜5).
 Thirdly, to apply APR analysis to liveness verification, we introduce a novel stronger validity of APR problems, called total validity, which requires not only finite but also infinite execution paths to reach target sets (Section˜6.1).
-Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to liveness verification (Section˜6.2).
-All omitted proofs of the claims are provided in both the appendix.
+Finally, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to liveness verification (Section˜6.2).
+All omitted proofs of the claims are provided in the appendix.
 As mentioned above, partial validity takes into account finite execution paths only.
@@
 Since “demonical validity” [CL18] (partial validity in this paper) does not consider infinite execution paths and we consider a stronger validity for all execution paths, we renamed “demonical validity” to “partial validity”, which originates from partial and total correctness of programs.
-Notice that total correctness of programs requires termination of programs, but total validity does not require termination but requires all (possibly infinite) execution paths reach elements in target sets.
+Notice that total correctness of programs requires termination of programs, but total validity does not require termination but requires all (possibly infinite) execution paths to reach elements in target sets.
 The main contributions of this paper are (1) a reformulation of the APR framework for ARSs, (2) a novel notion of validity—total validity—for APR predicates, and (3) a sufficient condition for total validity for, e.g., APR-based verification of liveness properties.
@@
 We say that $\tau$ is an execution path (of $\mathcal{A}$) if $\tau$ is maximal, i.e., $\tau$ is either finite ending with an irreducible element or infinite.
-Note that execution paths are defined coinductively by using two rules in $\mathsf{DVP}$ w.r.t. $\mathcal{A}$ [CL18] (Figure˜1 in this paper).
+Note that execution paths are defined coinductively by using two inference rules in $\mathsf{DVP}$ w.r.t. $\mathcal{A}$ [CL18] (Figure˜1 in this paper).
 Next, we define APR predicates over $A$.
@@
 We denote the greatest fixed point of $\widehat{\mathit{Rules}}$ by $\nu\widehat{\mathit{Rules}}$, i.e., $\nu\widehat{\mathit{Rules}}$ is the greatest set $Y$ ($\subseteq\mathcal{O}$) such that $Y\supseteq\widehat{\mathit{Rules}}(Y)$.
-An APR predicate $P\Rightarrow Q$ over $A$ is said to be partially valid (demonically valid [CL18]) w.r.t.
+An APR predicate $P\Rightarrow Q$ over $A$ is said to be partially valid (or demonically valid [CL18]) w.r.t.
 $\mathcal{A}$ if each execution path starting from an element in $P$ either eventually reaches an element in $Q$ or is infinite.
@@
 Definition 2.3 (partial validity [CL18]).
-An APR predicate $P\Rightarrow Q$ over $\mathcal{A}$ is said to be partially valid w.r.t.
+An APR predicate $P\Rightarrow Q$ over $A$ is said to be partially valid w.r.t.
 $\mathcal{A}$, written as $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$, if $(P\Rightarrow Q)\in\nu\widehat{\mathsf{DVP}(\mathcal{A})}$, where $\mathsf{DVP}(\mathcal{A})$ consists of the rules Subsumption and Step in Figure˜1.
@@
 $\mathcal{A}$.
-Regarding partial validity, APR predicates have the following properties.
+By definition, APR predicates trivially have the following properties for partial validity.
 Proposition 2.4.
@@
     $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow\emptyset$, if and only if any element in $P$ has no normal form, i.e., $\{t\in\mathit{NF\!}_{\mathcal{A}}\mid\exists s\in P.\ s\mathrel{\to_{\mathcal{A}}^{*}}t\}=\emptyset$.
-Proof 2.5.
-Trivial by definition.
-For an APR predicate $P\Rightarrow Q$, regarding partial validity, Proposition˜2.4 (1 and 2) implies soundness and completeness of splitting $P$ to $n$ sets $P_{1},\ldots,P_{n}$ (i.e., $P=P_{1}\cup\cdots\cup P_{n}$):
-  • $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$ if and only if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}(P_{1}\cup\cdots\cup P_{n})\Rightarrow Q$ (Proposition˜2.4-12).
-On the other hand, splitting the target set $Q$ to $Q_{1},\ldots,Q_{n}$ (i.e., $Q=Q_{1}\cup\cdots\cup Q_{n}$) is sound but not complete in general:
-  • if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$, then, $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow(Q_{1}\cup\cdots\cup Q_{n})$ (Proposition˜2.4-1), but
-  • the other direction does not hold in general.
+For an APR predicate $P\Rightarrow Q$, regarding partial validity, Proposition˜2.4 12 implies soundness and completeness of splitting $P$ to $n$ sets $P_{1},\ldots,P_{n}$ (i.e., $P=P_{1}\cup\cdots\cup P_{n}$): $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$ if and only if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}(P_{1}\cup\cdots\cup P_{n})\Rightarrow Q$.
+On the other hand, splitting the target set $Q$ to $Q_{1},\ldots,Q_{n}$ (i.e., $Q=Q_{1}\cup\cdots\cup Q_{n}$) is sound by Proposition˜2.4 1, but not complete in general: if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q_{i}$ for all $1\leq i\leq n$, then, $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow(Q_{1}\cup\cdots\cup Q_{n})$, but the other direction does not hold in general.
 The inference rules in $\mathsf{DVP}(\mathcal{A})$ can be considered a proof system for partial validity of APR predicates, i.e., $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ if and only if there exists a (possibly infinite) proof tree obtained from $P\Rightarrow Q$ by applying the rules of $\mathsf{DVP}(\mathcal{A})$.
@@
 $\{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c}\}$  (Step) $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$
-3 Reformulation of Inference Rules fo ARSs
-In this section, we reformulate the proof system $\mathsf{DVP}$ for partial validity w.r.t.
-ARSs in [CL18] to establish a one-to-one correspondence between the proof rules for partial validity w.r.t.
-ARSs and LCTRSs.
-By abstracting the inference rules in Figure˜2, we obtain inference rules for ARSs, which have one-to-one correspondence with the rules for LCTRSs.
+3 Reformulation of Inference Rules for ARSs
+In this section, we reformulate the proof system $\mathsf{DVP}$ for partial validity w.r.t. ARSs to establish a one-to-one correspondence between the inference rules for ARSs and LCTRSs.
+By abstracting the inference rules for LCTRSs in Figure˜2, we obtain inference rules for ARSs, which have one-to-one correspondence with the inference rules for LCTRSs.
 Definition 3.1 ($\mathsf{DVP}_{+}$).
-We define $\mathsf{DVP}_{+}(\mathcal{A})$ consisting of the rules in Figure˜3.
+We define $\mathsf{DVP}_{+}(\mathcal{A})$ consisting of the inference rules in Figure˜3.
 In applying Subs and Der of $\mathsf{DVP}_{+}(\mathcal{A})$ to $P\Rightarrow Q$, we can freely split the results of $P\setminus Q$ and $\partial_{\mathcal{A}}(P)$ to $n$ sets $P_{1},\ldots,P_{n}$, respectively.
@@
 Note that for an $\mathcal{A}$-runnable set $P$, we have that $\partial_{\mathcal{A}}(P)\neq\emptyset$ for Der, because $P\setminus\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$.
-Since constrained rewrite rules of an LCTRS may overlap, we allow $P_{i}$ and $P_{j}$ ($i\neq j$) to have overlaps, i.e., $P_{i}\cap P_{j}\neq\emptyset$.
+Note also that $P_{i}$ and $P_{j}$ ($i\neq j$) may overlap, i.e., $P_{i}\cap P_{j}\neq\emptyset$.
 $\mbox{\rm{\rm{Axiom}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P=\emptyset$.}\qquad\mbox{\rm{\rm{Subs}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ P_{1}\Rightarrow Q\quad\ldots\quad P_{n}\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\cap Q\neq\emptyset$,}$
@@
 Let $P,Q\subseteq A$. Then, at most one rule in $\mathsf{DVP}_{+}(\mathcal{A})$ is applicable to the APR predicate $P\Rightarrow Q$.
-Proof 3.3 (Proof (Sketch)).
-The side conditions of rules in $\mathsf{DVP}_{+}(\mathcal{A})$ are non-overlapping.
 For any ARS, $\mathsf{DVP}$ and $\mathsf{DVP}_{+}$ coincide.
@@
 For any ARS $\mathcal{A}$, $\nu\widehat{\mathsf{DVP}(\mathcal{A})}=\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$.
-Proof 3.5 (Proof (Sketch)).
-The application of Subsumption in $\mathsf{DVP}(\mathcal{A})$ can be simulated by Subs and Axiom in $\mathsf{DVP}_{+}(\mathcal{A})$ (with $n=1$).
-The application of Step in $\mathsf{DVP}(\mathcal{A})$ can be simulated by Subs and Der in $\mathsf{DVP}_{+}(\mathcal{A})$ (with $n=1$).
-Therefore, we have that $\nu\widehat{\mathsf{DVP}(\mathcal{A})}\subseteq\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$.
-The application of Axiom in $\mathsf{DVP}_{+}(\mathcal{A})$ can be simulated by Subsumption in $\mathsf{DVP}(\mathcal{A})$.
-The application of Subs in $\mathsf{DVP}_{+}(\mathcal{A})$ is followed by either Axiom or Der in $\mathsf{DVP}_{+}(\mathcal{A})$.
-The sequential application of Subs and Axiom can be simulated by Subsumption in $\mathsf{DVP}(\mathcal{A})$.
-The sequential application of Subs and Der can be simulated by Step in $\mathsf{DVP}(\mathcal{A})$.
-The application of Der in $\mathsf{DVP}_{+}(\mathcal{A})$—a special case of the successive application of Subs and Der—can be simulated by Step in $\mathsf{DVP}(\mathcal{A})$.
 The disproof criterion for LCTRSs [KN24jip] is formulated for ARSs as follows.
@@
 Let $P,Q\subseteq A$. If $P\cap Q=\emptyset$ and $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$, then $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$.
-Proof 3.7.
+Proof 3.5.
 We proceed by contradiction.
 Assume that $P\cap Q=\emptyset$, $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$, and $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$.
-Then, $P$ includes a normal form $s$ of $\mathcal{A}$: $s\in P\cap\mathit{NF\!}_{\mathcal{A}}$.
-Since $P\cap Q=\emptyset$, we have that $s\notin Q$.
-Since $s$ is a normal form of $\mathcal{A}$, $s$ itself is an execution path of $\mathcal{A}$, which starts from $P$.
+Then, there exists a normal form $s\in P\cap\mathit{NF\!}_{\mathcal{A}}$ such that $s\notin Q$.
 It follows from $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ that $s\in Q$.
 This contradicts the fact that $s\notin Q$.
-Proposition˜3.6 implies an inference rule for disproof (cf. [KN24jip]).
+Proposition˜3.4 implies an inference rule for disproof (cf. [KN24jip]).
 Definition 3.6 ($\mathsf{DVP}_{\bot}$).
-The set $\mathsf{DVP}_{\bot}(\mathcal{A})$ of inference rules consists of the rules in $\mathsf{DVP}_{+}(\mathcal{A})$ and the following rule for disproof:
+We define $\mathsf{DVP}_{\bot}(\mathcal{A})$ consisting of the inference rules in $\mathsf{DVP}_{+}(\mathcal{A})$ and the following rule for disproof:
 $\mbox{\rm{\rm{Dis}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ \bot\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }\penalty 10000\ \mbox{if $P\cap Q=\emptyset$, $P\neq\emptyset$, and $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$ }$
 Note that $\bot$ is used as a special case—the invalid one—of APR predicates.
-By definition, it is clear that for each APR predicate $P\Rightarrow Q$, exactly one rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ is applicable (cf. [KN24jip]).
+By definition, for each APR predicate, exactly one rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ is applicable (cf. [KN24jip]).
 Proposition 3.7.
 Let $P,Q\subseteq A$. Then, exactly one rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ is applicable to the APR predicate $P\Rightarrow Q$.
-Proof 3.10 (Proof (Sketch)).
-The side conditions of rules in $\mathsf{DVP}_{\bot}(\mathcal{A})$ are non-overlapping.
 4 Cyclic Proof System for APR Problems of ARSs
@@
 LCTRSs to the abstract APR framework reformulated in Section˜3.
-To this end, we first revisit rule Circ in Figure˜2, and then define APR pre-proofs for partial validity w.r.t.
+To this end, we first revisit Circ in Figure˜2, and then define APR pre-proofs for partial validity w.r.t.
 ARSs.
@@
 $\mbox{\rm{\rm{Cut}}}\penalty 10000\ \,\displaystyle\frac{\penalty 10000\ P\Rightarrow Q^{\prime}\qquad Q^{\prime}\Rightarrow Q\penalty 10000\ }{\penalty 10000\ P\Rightarrow Q\penalty 10000\ }$
-As in other proof systems, Cut must be powerful but needs a heuristic which may not be easy to automate.
-As a first step, we do not consider Cut and leave the introduction of Cut to the proof system as future work.
+As in other proof systems, Cut must be powerful but needs a heuristic for automation. As a first step, we leave the introduction of Cut to the proof system below as future work.
 In [KN24jip], a simpler proof system for partial validity w.r.t.
@@
 In the proof system, the circularity rule is not explicitly used, but the bud-companion relationship is used instead in proof trees.
-We reformulate the cyclic-proof system for partial validity w.r.t.
+We formulate the cyclic-proof system for partial validity w.r.t.
 ARSs.
@@
   • p is a partial mapping from $V$ to $V^{*}$ (we write $\mbox{\sl p}_{j}(v)$ for the $j$-th component of $\mbox{\sl p}(v)$) which is the $j$-th child of $v$, and
-  • for all nodes $v\in V$, $\mbox{\sl p}_{j}(v)$ is defined just in case $\mbox{\sl r}(v)$ is a rule with $k$ premises ($1\leq j\leq k)$, and $\frac{\mbox{\sl a}(\mbox{\sl p}_{1}(v))\penalty 10000\ \ldots\penalty 10000\ \mbox{\sl a}(\mbox{\sl p}_{k}(v))}{\mbox{\sl a}(v)}$ is an instance of rule $\mbox{\sl r}(v)$.
-[^3] If $k=0$, then $\mbox{\sl p}(v)=\epsilon$.
+  • for all nodes $v\in V$, $\mbox{\sl p}_{j}(v)$ is defined just in case $\mbox{\sl r}(v)$ is a rule with $k$ premises ($1\leq j\leq k)$, and $\frac{\mbox{\sl a}(\mbox{\sl p}_{1}(v))\penalty 10000\ \ldots\penalty 10000\ \mbox{\sl a}(\mbox{\sl p}_{k}(v))}{\mbox{\sl a}(v)}$ is an instance of rule $\mbox{\sl r}(v)$, where if $k=0$, then $\mbox{\sl p}(v)=\epsilon$.
 Note that a node $v\in V$ is a leaf if and only if either $\mbox{\sl p}(v)=\epsilon$ or $\mbox{\sl p}(v)$ is undefined.
@@
 An APR pre-proof for an APR predicate $P\Rightarrow Q$ w.r.t.
-$\mathcal{A}$ is a pair $(\mathcal{T},\xi)$ of an APR derivation tree $\mathcal{T}=(V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p})$ (with $v_{0}$ the root node) and a partial mapping $\xi$ from $V_{\mathit{open}}$ to $V\setminus V_{\mathit{open}}$ such that $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$, and for any open leaf $v$, if $\xi(v)$ is defined, then $\xi(v)$ is a node of $\mathcal{T}$ such that $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$ [^4] From the viewpoint of Cyc, the condition $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$” can be relaxed to $P\subseteq P^{\prime}$”, where $\mbox{\sl a}(v)=(P\Rightarrow Q)$ and $\mbox{\sl a}(v^{\prime})=(P^{\prime}\Rightarrow Q)$.
+$\mathcal{A}$ is a pair $(\mathcal{T},\xi)$ of an APR derivation tree $\mathcal{T}=(V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p})$ (with $v_{0}$ the root node) and a partial mapping $\xi$ from $V_{\mathit{open}}$ to $V\setminus V_{\mathit{open}}$ such that $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$, and for any open leaf $v$, if $\xi(v)$ is defined, then $\xi(v)$ is a node of $\mathcal{T}$ such that $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$ [^3] From the viewpoint of Cyc, the condition $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$” can be relaxed to $\mbox{\sl a}_{L}(v)\subseteq\mbox{\sl a}_{L}(\xi(v))$”.
 However, as a starting point, we do not use the relaxed condition, and leave it as future work.
@@
 Example 4.3.
-Let us consider the ARS $\mathcal{A}_{1}$ in Example˜2.6 again. Let $v_{0},v_{1},\ldots,v_{5}$ nodes such that
+Let us consider the ARS $\mathcal{A}_{1}$ in Example˜2.5 again. Let $v_{0},v_{1},\ldots,v_{5}$ nodes such that
   • $\mbox{\sl a}(v_{0})=(\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\})$, $\mbox{\sl r}(v_{0})=\mbox{\rm{\rm{Der}}}$, $\mbox{\sl p}(v_{0})=v_{1}$,
@@
 Then, $((\{v_{0},v_{1},\ldots,v_{5}\},\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is an APR pre-proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$ w.r.t.
-$\mathcal{A}_{1}$.
-The APR pre-proof is visualized in Figure˜4 (a), where $\dagger$} indicates the bud-companion relationship, i.e., $\xi(v_{3})=v_{0}$.
-$\dagger$} $v_{3}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$            (Axiom) $v_{5}:\penalty 10000\ \emptyset\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Subs) $v_{4}:\penalty 10000\ \{\mathsf{c}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Der)        $v_{2}:\penalty 10000\ \{\mathsf{b}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Subs)      $v_{1}:\penalty 10000\ \{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Der)      $\dagger$} $v_{0}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}{\dagger}$
+$\mathcal{A}_{1}$, which is visualized in Figure˜4 (a).
+$\dagger$} $v_{3}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$            (Axiom) $v_{5}:\penalty 10000\ \emptyset\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Subs) $v_{4}:\penalty 10000\ \{\mathsf{c}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Der)        $v_{2}:\penalty 10000\ \{\mathsf{b}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Subs)      $v_{1}:\penalty 10000\ \{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Der)       $\dagger$} $v_{0}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$
 ${v_{5}}$${v_{4}}$${v_{2}}$${v_{1}}$${v_{0}}$
-(a) APR pre-proof (b) the proof graph of (a)
-Figure 4: An APR pre-proof and its proof graph for the APR predicate $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$.
+(a) APR pre-proof (b) proof graph
+Figure 4: An APR pre-proof and its proof graph for the APR predicate $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$, where $\dagger$} indicates the bud-companion relationship.
 The proof graph obtained from a closed APR pre-proof is defined as follows.
@@
 Example 4.5.
-The proof graph of $((\{v_{0},v_{1},\ldots,v_{5}\},\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ in Example˜4.3 (i.e., Figure˜8 (a)) is illustrated in Figure˜8 (b).
-By definition, nodes of proof graphs have the following properties.
+The proof graph of $((\{v_{0},v_{1},\ldots,v_{5}\},\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ in Example˜4.3 (i.e., Figure˜4 (a)) is illustrated in Figure˜4 (b).
+By definition, nodes of proof graphs trivially have the following properties.
 Proposition 4.6.
-Let $(\mathcal{T},\xi)$ be a closed APR pre-proof, $(V,E)$ the proof graph of $(\mathcal{T},\xi)$, $v,v^{\prime}\in V$, $X\in\{\mbox{\rm{\rm{Axiom}}},\mbox{\rm{\rm{Subs}}},\mbox{\rm{\rm{Der}}},\mbox{\rm{\rm{Dis}}}\}$.
+Let $(\mathcal{T},\xi)$ be a closed APR pre-proof, $(V,E)$ the proof graph of $(\mathcal{T},\xi)$, $v,v^{\prime}\in V$, and $X\in\{\mbox{\rm{\rm{Axiom}}},\mbox{\rm{\rm{Subs}}},\mbox{\rm{\rm{Der}}},\mbox{\rm{\rm{Dis}}}\}$.
 Then, all of the following statements hold:
@@
     if $\mbox{\sl r}(v)=\mbox{\rm{\rm{Axiom}}}$, then $\mbox{\sl a}_{L}(v)=\emptyset$ and $v$ has no outedge.
-Proof 4.7.
-Trivial by definition.
 Cyclic proofs satisfy the global trace condition [Bro05].
@@
 In our case, such edges correspond to those from nodes with Der.
-To make proof graphs implicitly satisfy the global trace condition, we require companions to be nodes with Der; the form of APR predicates can be considered the same as that of sequents, and each APR predicate has exactly one constrained term corresponding to an atomic formula of premises of sequents; viewed in this light, traces correspond to paths of proof graphs of closed APR pre-proofs; every cycle goes through companions infinitely often, and thus every infinite trace follows an infinite path that goes through infinitely many nodes with Der.
+To make proof graphs implicitly satisfy the global trace condition, we require companions to be nodes with Der; the form of APR predicates can be considered the same as that of sequents, and each APR predicate has exactly one set corresponding to an atomic formula of premises of sequents; viewed in this light, traces correspond to paths of proof graphs of closed APR pre-proofs; every cycle goes through companions infinitely often, and thus every infinite trace follows an infinite path that goes through infinitely many nodes with Der.
 We now define APR (dis)proofs, which are APR pre-proofs satisfying certain conditions.
@@
    $\bot$  (Dis) $\{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c}\}$  (Der) $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$
-Figure 5: An APR pre-proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$, which is an APR disproof
+Figure 5: An APR pre-proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$, which is an APR disproof.
 APR proofs and disproofs are sound.
@@
     if there exists an APR disproof of $P\Rightarrow Q$, then $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$.
-Proof 4.11 (Proof (Sketch)).
-The first claim is proved by contradiction, analogously to the proof in [CL18, Theorem 3].
-For the second claim, the APR disproof for $P\Rightarrow Q$ has a path from the root to a leaf with $\bot$, which implies a finite execution path that starts from $P$ and does not include any element in $Q$.
 When $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, the APR predicate $P\Rightarrow Q$ has either an APR proof or an APR disproof.
@@
   2. 2.
-    if $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, and the emptiness, intersection emptiness, and $\mathcal{A}$-runnability problems for $A$ are decidable, then partial validity of $P\Rightarrow Q$ w.r.t.
-$\mathcal{A}$ is decidable.
-Proof 4.13 (Proof (Sketch)).
-For partial validity of $P\Rightarrow Q$, it suffices to consider APR predicates $P^{\prime}\Rightarrow Q$ such that $P^{\prime}\subseteq\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$.
-Since $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, there are only finitely many APR predicates to be considered, and thus we can construct a closed APR pre-proof by making generated nodes buds as much as possible.
-Since the closed APR pre-proof is either an APR proof or disproof, the first claim holds.
-For the second claim, since the side conditions of rules in $\mathsf{DVP}_{\bot}(\mathcal{A})$ are decidable by the assumption, we can construct a closed APR pre-proof, which is either an APR proof or disproof.
-Therefore, the partial validity of $P\Rightarrow Q$ w.r.t.
+    if $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite and the emptiness, intersection emptiness, and $\mathcal{A}$-runnability problems for $A$ are decidable, then partial validity of $P\Rightarrow Q$ w.r.t.
 $\mathcal{A}$ is decidable.
@@
 a system and a set $E$ of error states, we consider the property that any (possibly non-terminating) execution of the system never reaches any error state in $E$.
-This kind of safety properties is formulated for ARSs as follows:
+This kind of safety properties is formulated as a problem for ARSs as follows:
 Instance
@@
 Is there a (possibly infinite) execution path $s_{0}\to_{\mathcal{A}}s_{1}\to_{\mathcal{A}}\cdots$ such that $s_{0}\in P$ and $s_{i}\in E$ for some $i\geq 0$?
-This problem is the same as non-reachability from $P$ to $E$.
+This problem is the same as non-reachability from $P$ to $E$ w.r.t.
+$\mathcal{A}$.
 On the other hand, this problem can be reduced to the APR problem [KN23jlamp].
@@
   4. 4.
-    the non-reachability problem w.r.t.
-the error states $E$ is reduced to the APR predicate $I\Rightarrow\{\mathsf{success}\}$, where $I$ is the set of ground terms representing initial states of a system modeled by $\mathcal{R}$.
+    the non-reachability problem w.r.t. the error states $E$ is reduced to the APR predicate $I\Rightarrow\{\mathsf{success}\}$, where $I$ is the set of ground terms for initial states of a system.
 Note that the introduced constants are normal forms of the modified LCTRS.
@@
 Usually, error states are irreducible, and thus we consider only irreducible error states.
-If an error state to be considered is reducible, then, as in [KN23jlamp], we introduce a fresh constant such as $\mathsf{error}$ above, together with constrained rewrite rules from the error states to $\mathsf{error}$, which is considered a dummy error state in the modified system.
+If an error state is reducible, then, as in the approach above for LCTRSs, we introduce a fresh constant such as $\mathsf{error}$, together with the additional reduction from the error states to $\mathsf{error}$, which is considered a dummy error state in the modified system.
 Let us reconsider to reduce non-reachability (from $P$ to $E$ w.r.t.
@@
 Then, $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ if and only if there is no finite execution path of $\mathcal{A}$ that starts with an element in $P$ and includes an element in $E$.
-Proof 5.3.
-Proof (Sketch) Both the only-if and if parts are proved by contradiction.
+${\mathit{PG}_{0}:\penalty 10000\ \penalty 10000\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\mathsf{noncrit}_{0}$$\footnotesize{\langle b_{0},x\rangle:=\langle\mathsf{true},1\rangle}$$\mathsf{wait}_{0}$$\footnotesize{x=0\,\lor\,\lnot b_{1}}$$\mathsf{crit}_{0}$$\footnotesize{b_{0}:=\mathsf{false}}$
+${\mathit{PG}_{1}:\penalty 10000\ \penalty 10000\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\mathsf{noncrit}_{1}$$\footnotesize{\langle b_{1},x\rangle:=\langle\mathsf{true},0\rangle}$$\mathsf{wait}_{1}$$\footnotesize{x=1\,\lor\,\lnot b_{0}}$$\mathsf{crit}_{1}$$\footnotesize{b_{1}:=\mathsf{false}}$
+Figure 6: Program graph $\mathit{PG}_{i}$ ($i=0,1$) for Peterson’s mutual exclusion algorithm.
 Example 5.3.
-Let us consider the program graph in Figure˜6 for Peterson’s mutual exclusion algorithm [BK08, Example 2.25].
-Two processes $P_{0},P_{1}$ share Boolean variables $b_{0},b_{1}$ and an integer variable $x$; The Boolean variables $b_{i}$ indicates that $P_{i}$ wants to enter the critical section $\mathsf{crit}_{i}$; The integer variable $x$ stores the identifier of the process that has priority for the critical section at that time.
+Let us consider the program graphs $\mathit{PG}_{0},\mathit{PG}_{1}$ in Figure˜6 for Peterson’s mutual exclusion algorithm [BK08, Example 2.25].
+Two processes $P_{0},P_{1}$ specified by $\mathit{PG}_{0},\mathit{PG}_{1}$, respectively, share Boolean variables $b_{0},b_{1}$ and an integer variable $x$; The Boolean variables $b_{i}$ indicates that $P_{i}$ wants to enter the critical section $\mathsf{crit}_{i}$; The integer variable $x$ stores the identifier of the process that has priority for the critical section at that time.
 A state of the asynchronous integer transition systems (AITS, for short) with shared variables [BK08] consisting of $\mathit{PG}_{0}$ and $\mathit{PG}_{1}$ are a tuple of a location of $\mathit{PG}_{0}$, a location of $\mathit{PG}_{1}$, and an assignment from variables $b_{0},b_{1},x$ to values.
 The initial states of the AITS are $\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle$ with $m\in\{0,1\}$.
-Let $\mathit{State}_{[}i]$ be the set $\{\mathsf{noncrit}_{i},\mathsf{wait}_{i},\mathsf{crit}_{i}\}$ ($i\in\{0,1\}$), $\mathbb{B}$ the set $\{\mathsf{true},\mathsf{false}\}$, and $\mathcal{S}$ the set of assignments for variables $b_{0},b_{1}:\mbox{\sl{bool}}$ and $x:\mbox{\sl{int}}$: $\mathcal{S}=\{\{b_{0}\mapsto v_{0},b_{1}\mapsto v_{1},x\mapsto m\}\mid v_{0},v_{1}\in\mathbb{B},\penalty 10000\ m\in\mathbb{Z}\}$.
+Let $\mathit{State}_{i}$ be the set $\{\mathsf{noncrit}_{i},\mathsf{wait}_{i},\mathsf{crit}_{i}\}$ ($i\in\{0,1\}$), $\mathbb{B}$ the set $\{\mathsf{true},\mathsf{false}\}$, and $\mathcal{S}$ the set of assignments for variables $b_{0},b_{1}:\mbox{\sl{bool}}$ and $x:\mbox{\sl{int}}$: $\mathcal{S}=\{\{b_{0}\mapsto v_{0},b_{1}\mapsto v_{1},x\mapsto m\}\mid v_{0},v_{1}\in\mathbb{B},\penalty 10000\ m\in\mathbb{Z}\}$.
 We denote a state $\langle\,p_{0},p_{1},\sigma\,\rangle$ by $\langle\,p_{0},p_{1},\sigma(b_{0}),\sigma(b_{1}),\sigma(x)\,\rangle$.
-The AITS is represented by the following ARS:
+The AITS consisting of $P_{0}$ and $P_{1}$ is represented by the following ARS:
 $\mathcal{A}_{2}=(\{\langle\,p_{0},p_{1},\sigma\,\rangle\mid p_{0}\in\mathit{State}_{0},\,p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\},\to_{\mathcal{A}_{2}})$
@@
 Since all the error states are reducible w.r.t.
-$\mathcal{A}_{2}$, we introduce a fresh element $\mathsf{error}$ as an error state to be verified, and add the reduction from the error states to $\mathsf{error}$: $\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle\to_{\mathcal{A}_{2}}\mathsf{error}$, where $\sigma\in\mathcal{S}$.
+$\mathcal{A}_{2}$, we introduce a fresh element $\mathsf{error}$ as a dummy error state to be verified, and add the reduction from the original error states to $\mathsf{error}$: $\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle\to_{\mathcal{A}_{2}}\mathsf{error}$, where $\sigma\in\mathcal{S}$.
 We let $\mathcal{A}_{2}^{\prime}$ be the modified ARS $(\{\langle\,p_{0},p_{1},\sigma\,\rangle\mid p_{0}\in\mathit{State}_{0},\,p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\},\to_{\mathcal{A}_{2}^{\prime}})$, where ${\to_{\mathcal{A}_{2}^{\prime}}}={\to_{\mathcal{A}_{2}}}\cup\{(\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle,\mathsf{error})\mid\sigma\in\mathcal{S}\}$.
-The race freedom is reduced to the following APR predicate:
-$\{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\emptyset$
-We obtain an APR proof illustrated in Figure˜7, and thus, by Theorem 4.10, this APR predicate is partially valid w.r.t.
+The race freedom is reduced to the APR predicate $(1)\penalty 10000\ \{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\emptyset$.
+We obtain an APR proof illustrated in Figure˜7, and thus, by Theorem 4.9, the APR predicate (1) is partially valid w.r.t.
 $\mathcal{A}_{2}^{\prime}$.
 Therefore, by Theorem˜5.2, the AITS defined by the program graphs in Figure˜6 is race free and, by Proposition˜2.4 5, there is no normal form reachable from any initial state.
-${\mathit{PG}_{0}:\penalty 10000\ \penalty 10000\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\mathsf{noncrit}_{0}$$\footnotesize{\langle b_{0},x\rangle:=\langle\mathsf{true},1\rangle}$$\mathsf{wait}_{0}$$\footnotesize{x=0\,\lor\,\lnot b_{1}}$$\mathsf{crit}_{0}$$\footnotesize{b_{0}:=\mathsf{false}}$
-${\mathit{PG}_{1}:\penalty 10000\ \penalty 10000\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\mathsf{noncrit}_{1}$$\footnotesize{\langle b_{1},x\rangle:=\langle\mathsf{true},0\rangle}$$\mathsf{wait}_{1}$$\footnotesize{x=1\,\lor\,\lnot b_{0}}$$\mathsf{crit}_{1}$$\footnotesize{b_{1}:=\mathsf{false}}$
-Figure 6: Program graph $\mathit{PG}_{i}$ ($i=0,1$) for Peterson’s mutual exclusion algorithm.
 (8)    (3)  (Der) (9)  (Der)    (4)     (9)  (Der) (5)     (Der)       (2)     (2)  (Der) (10)  (Der) (6)     (10)    (2)   (3)     (11)  (Der)    (7)  (Der)      (3)  (Der) (1) $\{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\emptyset$
@@
 Then, $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ if and only if $(A\cup\{\mathsf{any}\},{\to_{\mathcal{A}}}\cup\{(s,\mathsf{any})\mid s\in A\setminus E\})\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow\{\mathsf{any}\}$.
-Proof 5.6 (Proof (Sketch)).
-Both the only-if and if parts are proved by contradiction.
-One may think that non-reachability analysis is more suitable for verification of safety properties.
-However, APR analysis is essentially very similar to non-reachability analysis.
-For non-reachability from $P$ to $E$, all states reachable from $P$ are verified, unless an error state reachable from $P$ is detected.
-On the other hand, during the construction of an APR pre-proof, all states reachable from $P$ are verified, unless a subset $P^{\prime}$ with $P^{\prime}\cap E\neq\emptyset$ is detected.
+One may think that non-reachability analysis is more suitable for safety properties.
+However, APR analysis is essentially very similar to non-reachability analysis; for non-reachability from $P$ to $E$, all states reachable from $P$ are examined, unless an error state reachable from $P$ is detected; during the construction of an APR pre-proof, all states reachable from $P$ are examined, unless a subset $P^{\prime}$ with $P^{\prime}\cap E\neq\emptyset$ is detected.
 6 Reduction of Liveness Properties to APR Problems
@@
 a system, a source set $P$, and a set $Q$ for “something good”, we consider the property that any (possibly non-terminating) execution of the system starting from any state in $P$ reaches a state in $Q$.
-This kind of safety properties is formulated for ARSs as follows:
+This kind of safety properties is formulated as a problem for ARSs as follows:
 Instance
@@
 In this section, we introduce a stronger validity, called total validity, which takes into account all (possibly infinite) execution paths.
-Then, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to verification of liveness properties.
+Then, for a partially valid APR predicate with an APR proof, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to verification of liveness properties.
 6.1 Total Validity of APR Predicates w.r.t. ARSs
@@
 Theorem 6.3.
-Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be a partially valid APR proof of an APR predicate $P\Rightarrow Q$ w.r.t.
+Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be a partially valid APR proof for an APR predicate $P\Rightarrow Q$ w.r.t.
 $\mathcal{A}$, and $(V^{\prime},E)$ the proof graph of $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$.
 Then, $(V^{\prime},E)$ is acyclic if and only if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.
-Proof 6.4 (Proof (Sketch)).
-Both the only-if and if parts are proved by contradiction.
-For the if part, we consider a bud node $v$ and its companion $v^{\prime}$ in a cycle such that $\xi(v)=v^{\prime}$.
-By the assumption, we have that $\mbox{\sl a}_{L}(v)=\mbox{\sl a}_{L}(v^{\prime})$.
-It follows from Proposition˜4.6 23 that for any element $s\in\mbox{\sl a}_{L}(v)$, there exists an element $s^{\prime}\in\mbox{\sl a}_{L}(v^{\prime})=\mbox{\sl a}_{L}(v)$ such that $s^{\prime}\to_{\mathcal{A}}^{+}s$ and the reduction has no element in $Q$.
-Using this claim, we can show an infinite execution path that starts with an element in $P$ and does not include any element in $Q$.
-This contradicts the assumption that $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.
-We call a closed APR pre-proof acyclic if the proof graph of the APR pre-proof is acyclic.
+Once we construct an APR proof, the total validity is decidable.
+Corollary 6.4.
+Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be an APR proof for a partially valid APR predicate $P\Rightarrow Q$ w.r.t.
+$\mathcal{A}$.
+Then, total validity of $P\Rightarrow Q$ w.r.t.
+$\mathcal{A}$ is decidable.
+Proof 6.5.
+It is decidable whether a finite graph is acyclic, and thus, by Theorem˜6.3, totally validity is decidable.
+We call a closed APR pre-proof acyclic if its proof graph is acyclic.
 As a consequence of Theorem˜6.3, for an APR predicate $P\Rightarrow Q$, the existence of an acyclic APR pre-proof is a sufficient condition for total validity of $P\Rightarrow Q$ w.r.t.
@@
 $\mathcal{A}$, then $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.
-Once we construct an APR proof, the total validity is decidable.
-Corollary 6.6.
-Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be an APR proof of a partially valid APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$. Then, both of the following statements hold:
-  1. 1.
-    $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$ if and only if the proof graph of $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is acyclic, and
-  2. 2.
-    total validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
-Proof 6.7.
-The first claim follows Theorem˜6.3. It is decidable whether a finite graph is acyclic. Thus, the second claim holds.
 Example 6.7.
 Let us continue with Example˜6.2.
-We have an acyclic APR proof for the APR predicate (12).
-Therefore, by Corollary˜6.5, the APR predicate (12) is totally valid w.r.t.
+We have an acyclic APR proof for the APR predicate (12) shown in Figure˜8.
+Therefore, by Corollary˜6.6, the APR predicate (12) is totally valid w.r.t.
 $\mathcal{A}_{2}$, and thus the AITS is starvation free for $P_{0}$.
-    (Axiom) (17)  (Subs) (13)        (Axiom) (17)  (Subs) (18)  (Der) (14)†        (Axiom) (17)  (Subs) (21)   (14)†  (Der)     (19)      (Der)     (15)        (Axiom) (17)  (Subs) (20)   (14)†  (Der)     (16)  (Der)                  (12)
+         (Axiom) (17)  (Subs) (13)        (Axiom) (17)  (Subs) (18)  (Der) (14)†        (Axiom) (17)  (Subs) (21)   (14)†  (Der)     (19)      (Der)     (15)        (Axiom) (17)  (Subs) (20)   (14)†  (Der)     (16)  (Der) (12) $\{\langle\,\mathsf{wait}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\sigma\in\mathcal{S},\sigma(b_{0})\,{=}\,\mathsf{true}\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\sigma\in\mathcal{S}\}$
   • (13)
@@
 The work in this paper is the adaptation of the APR frameworks in [CL18, KN23jlamp] to ARSs, providing an abstract foundation for APR analysis.
-$\mathbb{K}$ framework [RS10] is a more general setting of rewriting than LCTRSs, and the proof system for all-path reachability has been implemented in $\mathbb{K}$ [SCMMSR14, SCMMSR19].
+The cyclic-proof system for partial validity of APR predicates in this paper is a simplified adaptation of the well-known cyclic-proof system $\mbox{CLKID}^{\omega}$ [Bro06phd, Chapter 5].
+Compared with sequents, APR predicates are very simple, and our APR proofs always satisfy the global trace condition.
+Runtime-error verification by means of APR analysis w.r.t.
+constrained rewrite systems such as LCTRSs has been investigated.
+$\mathbb{K}$ framework [RS10] is a more general setting of constrained rewriting than LCTRSs, and the proof system for partial validity of APR predicates consisting of constrained terms has been implemented in $\mathbb{K}$ [SCMMSR14, SCMMSR19].
 The race-freedom of Peterson’s mutual exclusion algorithm has been proved in [SCMMSR19] by means of the APR approach, while starvation freedom is not considered.
 A comparison of all-path reachability logic with CTL* can be seen in [SCMMSR19].
-Our previous work [KN23padl, KN24rp] for the verification of starvation freedom by means of APR analysis with LCTRSs reduces starvation freedom to safety APR predicates, and does not use the approach to liveness properties in this paper.
 Our previous APR-based approach in [KN23jlamp] to safety verification is to add rules to a given LCTRS so as to make any finite prefix of all (possible infinite) execution paths finite execution paths of the modified LCTRS.
@@
 Viewed in this light, in practice, these two approaches are similar, and the approach in this paper is an abstract framework of the previous one for LCTRSs.
-The cyclic-proof system for partial validity of APR predicates in this paper is a simplified adaptation of the well-known cyclic-proof sytem $\mbox{CLKID}^{\omega}$ [Bro06phd, Chapter 5].
-Compared with sequents, APR predicates are very simple, and our APR proofs always satisfy the global trace condition.
+Our previous work [KN23padl, KN24rp] for the verification of starvation freedom by means of APR analysis with LCTRSs reduces starvation freedom to safety APR predicates, and does not use the approach to liveness properties in this paper.
+On the other hand, the previous work [KN24rp] deals with process-fairness for starvation freedom.
 8 Conclusion
@@
 Finally, we introduced total validity of APR predicates w.r.t.
-ARSs and we showed if there is an acyclic APR proof for an ARP predicate w.r.t.
+ARSs and showed if there is an acyclic APR proof for an APR predicate w.r.t.
 an ARS, then the APR predicate is totally valid w.r.t.
@@
 Our APR pre-proofs do not consider the generalization for bud nodes.
-To be more precise, in the current definition, a bud and its companion have the same APR predicate.
-As described in Footnote˜4, the relaxation of bud-companion relationships—an APR predicate can be more general than that of its companion—is one of our future work.
+To be more precise, in the current definition, a bud node and its companion have the same APR predicate.
+As described in Footnote˜3, the relaxation of bud-companion relationships—an APR predicate can be more general than that of its companion—is one of our future work.
 A future direction of this research is to deal with the full version of the rule Circ, i.e., “cut” in APR pre-proofs.
@@
     • Case where $P\cap Q\neq\emptyset$. Rule Sub is applicable, but the others are not.
-See 3.4
+Therefore, the claim holds.
+See 3.3
 Proof A.2.
@@
 The application of Der in $\mathsf{DVP}_{+}(\mathcal{A})$ is a special case of the successive application of Subs and Der in $\mathsf{DVP}(\mathcal{A})$.
-See 3.9
+See 3.7
 Proof A.3.
@@
 In this case, by definition, Dis is applicable.
-See 4.10
+See 4.9
 Proof A.4.
@@
 Therefore, we have that $\mathcal{A}\mathrel{\mathrel{\not\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$.
-See 4.12
+See 4.10
 Proof A.5.
 We first show a naive construction of APR pre-proofs of $P\Rightarrow Q$, where we do not split any source sets.
-It follows from Proposition˜3.9 that for each APR predicate, there is exactly one applicable rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$.
+It follows from Proposition˜3.7 that for each APR predicate, there is exactly one applicable rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$.
 Let $v_{0}$ be the root of a derivation tree we construct now, where $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$.
@@
 This contradicts the assumption that there is no execution path of $\mathcal{A}$ that starts with an element in $P$ and includes an element in $E$.
-See 5.5
+See 5.4
 Proof A.7.